### v1

# Psychiatry-Bench: A Multi-Task Benchmark for LLMs in Psychiatry

Abstract
Large language models (LLMs) hold great promise in enhancing psychiatric practice, from improving diagnostic accuracy to streamlining clinical documentation and therapeutic support. However, existing evaluation resources heavily rely on small clinical interview corpora, social media posts, or synthetic dialogues, which limits their clinical validity and fails to capture the full complexity of psychiatric reasoning. In this work, we introduce PsychiatryBench, a rigorously curated benchmark grounded exclusively in authoritative, expert-validated psychiatric textbooks and casebooks (e.g., DSM-5-TR Clinical Cases, Stahl’s Essential Psychopharmacology). PsychiatryBench comprises eleven distinct question-answering tasks ranging from diagnostic reasoning and treatment planning to longitudinal follow-up, management planning, clinical approach, sequential case analysis, and multiple-choice/extended matching formats totaling over 5,300 expert-annotated items. We evaluate a diverse set of frontier LLMs (including Google Gemini, DeepSeek, LLaMA 3, and QWQ-32) alongside leading open-source medical models (e.g., OpenBiloLLM, MedGemma) using both conventional metrics and an "LLM-as-judge" similarity scoring framework. Our results reveal substantial gaps in clinical consistency and safety, particularly in multi-turn follow-up and management tasks, underscoring the need for specialized model tuning and more robust evaluation paradigms. PsychiatryBench offers a modular, extensible platform for benchmarking and improving LLM performance in high-stakes mental health applications.
Keywords PsychiatryBench  $\cdot$ Large Language Models  $\cdot$ Psychiatric Benchmarking  $\cdot$ Diagnostic evaluation  $\cdot$ Mental Illness
1 Introduction
Large language models (LLMs), a subset of artificial intelligence (AI), have the potential to completely transform the way mental illness is identified, diagnosed, and treated. In 2019, mental disorders accounted for 5.1% of the world’s disease burden, with over 280 million people suffering from depressive disorders and around 703,000 people dying from suicide [1]. Between 2024 and 2040, the total economic burden of mental health problems in the United States is expected to reach around \$14 trillion, including direct medical expenses, emergency room visits, lost productivity, and premature death [2].
New avenues for mental health research were made possible with the introduction of Natural Language Processing (NLP). Early applications focused on analyzing text from self-report questionnaires and clinical interviews using techniques like the Linguistic Inquiry and Word Count (LIWC) [3]. A significant advancement came with transformer-based models such as BERT [4], which learn complex linguistic representations from large text corpora without manual feature engineering. These models have shown strong performance in tasks like sentiment analysis, suicide risk assessment, and detecting mental health disorders from social media content [5].
However, despite these advances, there remains a critical gap: existing evaluation frameworks rarely test models on tasks that reflect the depth and complexity of real-world psychiatric decision-making. Most benchmarks rely on simplified labels or non-clinical data, limiting their relevance to actual clinical practice. To address this, our study introduces a new benchmark grounded in expert-curated psychiatric sources and designed to assess nuanced reasoning tasks such as differential diagnosis and treatment planning. This framework not only raises the bar for clinical relevance in LLM evaluation but also enables robust comparison across state-of-the-art models.
The integration of LLMs into mental healthcare represents a significant technological shift, with the potential to transform clinical practice across the entire psychiatric workflow[6]. In diagnostics, LLMs demonstrate effectiveness in detecting disorders like depression from textual data; in therapeutic interventions, through conversational agents; and in clinical documentation, their automated notes have been preferred by therapists in blind tests over human-written ones[7]. Intriguingly, some studies suggest that LLMs can even surpass human physicians in expressing empathy in written communication[8]. This transformative potential, however, is counterbalanced by significant risks that necessitate an emerging field that can be termed ’Evaluation Psychiatry.’ A primary concern is the propensity of LLMs to generate inaccurate or nonsensical responses, known as "hallucinations," which raises substantial safety concerns in a clinical domain where errors can have profound consequences pandey2024harnessinglargelanguagemodels. This tension between the promise of democratizing mental health access and the ethical imperative of patient safety underscores the critical need for rigorous, clinically-grounded evaluation frameworks.
The root of this tension can be traced to the evolution of the datasets used to evaluate these models. Early, influential resources like the Distress Analysis Interview Corpus (DAIC-WOZ) provided clinically relevant transcripts that enabled foundational methods, but these datasets were small and faced challenges in generalizability [9]. To achieve greater scale, a second wave of datasets leveraged social media and public surveys, such as the subreddit Mental Health Posts dataset[10],[11]. Models like MentalBERT, pretrained on millions of sentences from mental health subreddits, achieved high accuracy (over 94%) on simplified tasks like binary depression classification[12]. While valuable for demonstrating potential, this approach has a critical flaw: its reliance on unverified, non-clinical data oversimplifies complex psychiatric disorders and risks reflecting trends in self-diagnosis rather than formal clinical presentations.
More recently, the field has begun to address this by developing more complex, conversational benchmarks. MentalChat16K, for example, combines synthetic data with anonymized clinical transcripts to evaluate an LLM’s ability to engage in nuanced dialogue[13]. Reflects a maturation of the field, moving beyond simple classification. Nonetheless, this third wave introduces new methodological challenges. The reliance on synthetic data risks creating benchmarks that do not capture the full complexity of real human interaction. At the same time, the use of LLM-as-a-judge evaluation methods introduces a circular dependency, where the technology being assessed is also the tool for its assessment, potentially masking its own biases and unreliability.
This analysis of the current landscape reveals a critical and persistent gap: the lack of a comprehensive benchmark that is simultaneously grounded in authoritative, expert-validated knowledge and designed to evaluate the full spectrum of complex clinical reasoning. There is a stark contrast between the datasets currently in use and the foundational knowledge sources that form the bedrock of clinical psychiatric practice. Authoritative textbooks and casebooks, such as the DSM-5-TR Clinical Cases[14] and Stahl’s Essential Psychopharmacology[15], represent knowledge curated and validated by experts over decades. These resources, which illustrate the complexities of differential diagnosis and treatment planning, stand as the gold standard against which an LLM’s clinical acumen and safety must ultimately be measured.
The present study is positioned to address these limitations directly. We introduce a novel evaluation framework meticulously constructed from the foundational psychiatric textbooks described above, deliberately prioritizing clinical accuracy, safety, and alignment with established medical practice. Our contribution is threefold:
  • We construct a novel, diverse dataset curated from authoritative clinical casebooks, diagnostic manuals, and specialized guides, ensuring that our evaluation is grounded in validated, expert knowledge rather than unverified or synthetic data.
  • We design eleven distinct and complex QA task types, ranging from diagnostic reasoning and treatment planning to sequential, case-based analysis, to create a framework that evaluates the sophisticated cognitive functions required in real-world psychiatry, moving far beyond simple classification.
  • We deploy this framework to conduct a rigorous evaluation of leading frontier models, establishing a crucial benchmark that provides novel insights into the current capabilities and inherent risks of LLMs in a clinical context.
  • A comparative evaluation of several state-of-the-art LLMs, including Gemini 2.5 Pro, Gemini 2.5 flash, Claude Opus, and Qwen/QWQ-32, selected for their strong performance across reasoning and safety benchmarks, and for their relevance in current clinical LLM discourse.
This work, therefore, moves beyond testing what LLMs can do on simplified tasks and begins to assess their readiness for the complex, rigorous, high-stakes reality of clinical mental healthcare. The remainder of the paper is structured as follows: Section 2 provides a detailed review of LLM applications in mental health and the landscape of psychiatric datasets. Section 3 describes the design, sourcing, and annotation of our novel psychiatric QA dataset. Section 4 outlines the experimental framework, and Section 5 details the implementation. Section 6 presents and interprets our results. Section 7 discusses study limitations and future work, and Section 8 concludes the paper by summarizing our key contributions. Finally, Ethical, Copyright, and Fair Use Statements are presented.
2 Related Work
In this section, we review the evolving landscape of LLMs in psychiatry and mental health, focusing on both their application potential and the underlying evaluation frameworks. We begin by surveying clinical uses of LLMs, ranging from diagnostics and therapy to documentation, and examining the ethical and safety concerns arising from their deployment. We then analyze the major datasets that have shaped model development in this space, highlighting limitations in clinical grounding, generalization, and task complexity. Finally, we discuss the importance of using expert-validated knowledge sources, such as psychiatric casebooks and diagnostic manuals, to develop robust, clinically meaningful benchmarks.
2.1 The Emergence of LLMs in Mental Healthcare
The integration of LLMs into healthcare represents a significant technological shift, with the potential to transform clinical practice, particularly in mental health [16]. The field is witnessing a rapid expansion of LLM applications across the entire psychiatric workflow. In diagnostics, LLMs demonstrate effectiveness in detecting mental health disorders and suicidal ideation from textual data, often sourced from social media, and they can translate complex diagnostic manuals like the ICD-11 into inspectable logic programs for expert review [17]. For therapeutic interventions, conversational agents such as Woebot and Wysa, built on principles of cognitive behavioral therapy, are providing accessible and de-stigmatized eHealth services. Intriguingly, some studies suggest that LLMs can even surpass human physicians in expressing empathy in written communication [18]. Furthermore, in clinical documentation, LLMs offer dramatic efficiency gains by automating the generation of medical notes. In this task, therapists have preferred their output in blind tests over human-written notes [19].
This transformative potential, however, is counterbalanced by significant risks that necessitate an emerging field that can be termed ’Evaluation Psychiatry.’ A primary concern is the propensity of LLMs to generate inaccurate or nonsensical responses, known as "hallucinations," which raises substantial safety concerns in a clinical domain where errors can have profound consequences [20]. This risk mandates stringent ethical guidelines and a "human-in-the-loop" paradigm where AI outputs are subject to expert validation. Clinicians also express ambivalence, worrying about over-reliance on "black box" algorithms, the erosion of clinical autonomy, and the potential for training data to embed biases related to socioeconomic status or geography, which could exacerbate healthcare disparities [21]. This tension between the promise of democratizing mental health access and the ethical imperative of patient safety underscores the critical need for rigorous, clinically-grounded evaluation frameworks.
2.2 The Landscape of Datasets for Psychiatric Evaluation
The evolution of LLM evaluation in mental health can be traced through the datasets used to train and test them. Early and influential resources include clinical interview corpora like the Distress Analysis Interview Corpus (DAIC-WOZ) and its extension, E-DAIC.[22] These datasets, containing transcripts of semi-structured interviews and corresponding psychiatric assessment scores (e.g., PHQ-8 for depression), have been instrumental. They enabled innovative methods such as the Language Model for Impersonation-based Questionnaire Completion (LMIQ) framework, which attempts to bridge unstructured patient narratives with quantitative metrics [18]. While a significant step, these methods face challenges in generalization beyond their specific benchmarks.
A second wave of datasets has leveraged the scale of social media and public surveys. Resources like the Reddit Mental Health Posts dataset, the Depression Detection Using Text corpus, and various mental health surveys have provided vast amounts of natural language data for tasks like sentiment analysis and binary classification [23]. Models like Mental BERT and Mental RoBERTa were pretrained on millions of sentences from mental health-related subreddits to improve performance in these tasks [17]. However, the clinical validity of social media data is a significant concern due to its unverified nature, oversimplification of symptoms, and the risk of reflecting trends in self-diagnosis rather than formal clinical presentations.
More recently, the field has begun to address the need for more complex, conversational benchmarks. MentalChat16K, for instance, is a benchmark for conversational mental health assistance that combines synthetic data with anonymized clinical transcripts. [13]. Simultaneously, researchers have also investigated automated approaches for enhancing dataset richness and annotation quality. For example, Hassan et al.(2024) introduced a synthetic labeling framework that transforms traditional single-disorder datasets into multi-label corpora by leveraging zero-shot and prompt-driven LLM annotation strategies. Their method, exemplified by the SPAADE-DR dataset, better captures co-occurring disorders such as depression, anxiety, PTSD, ADHD, and eating disorders, thus enabling more nuanced diagnostic modeling[24]. This reflects a maturation of the field, moving beyond simple classification towards evaluating LLMs on their ability to engage in nuanced, therapeutic dialogue. Nonetheless, the reliance on synthetic data and LLM-as-a-judge evaluation methods introduces new layers of potential bias and unreliability.
2.3 Foundational Knowledge Sources in Psychiatric Practice
To ensure clinical relevance and safety, any robust evaluation of LLMs in psychiatry must be grounded in the authoritative knowledge that defines the field. The gold standard for this knowledge is found in established psychiatric textbooks and clinical guides, which provide the basis for professional training and practice. The following sources represent the bedrock of clinical knowledge that LLMs must be tested against:
  • Diagnostic Manuals and Casebooks: Books like DSM-5 Clinical Cases and its successor, DSM-5-TR Clinical Cases, are essential companions to the official diagnostic manual. They bring diagnostic criteria to life through realistic narratives and detailed clinical discussions, illustrating the complexities of comorbidity and differential diagnosis. Similarly, texts such as 100 Cases in Psychiatry and Case Files Psychiatry provide high-yield clinical scenarios designed to sharpen practical decision-making skills, making them ideal for evaluating an LLM’s clinical reasoning.
  • Specialized Clinical Guides: Deep domain expertise is captured in specialized resources. Stahl’s Essential Psychopharmacology and its case-based companions are seminal texts on the neurobiological basis and practical application of psychiatric medications, crucial for evaluating treatment-related tasks. Likewise, Geriatric Psychiatry provides comprehensive knowledge on the unique diagnostic and treatment considerations for elderly populations.
  • Structured Knowledge and Assessment Tools: To test factual recall and application of diagnostic criteria in a standardized format, resources like DSM-5-TR Self-Exam Questions and MCQs in Psychology are invaluable. They provide thousands of multiple-choice questions that mirror the format of professional board examinations, offering a structured way to measure an LLM’s knowledge base.
These foundational texts, curated and validated by experts over decades, stand in stark contrast to the unverified and often transient nature of data from social media. Grounding an evaluation dataset in these sources is a deliberate choice to prioritize clinical accuracy, safety, and alignment with established medical practice.
2.4 Synthesis of Prior Work and Remaining Gaps
In reviewing the landscape, it is clear that prior work has made significant strides. LLMs have proven highly effective at classification tasks, such as achieving a balanced accuracy of 94.8% for depression detection and 96.2% for PTSD when incorporated via decision-level fusion in multimodal systems [25]. They have also demonstrated utility in administrative roles and have shown surprising capabilities in generating empathetic responses. These successes, however, also illuminate what is missing. The field’s reliance on datasets derived from social media, while useful for scale, lacks the clinical rigor required for high-stakes applications. Evaluation has often been limited to simple classification or sentiment analysis, failing to assess the complex, multi-step reasoning that defines clinical practice. Even with more advanced methods like LMIQ, challenges of generalization and the risk of hallucination persist.
This analysis reveals a critical gap in the literature: a comprehensive benchmark that is simultaneously grounded in authoritative, expert-validated knowledge and designed to evaluate the full spectrum of complex clinical reasoning. There is an urgent need to move beyond binary classification and simple QA to assess an LLM’s ability to perform tasks that require nuanced understanding, such as formulating a differential diagnosis, creating a multi-stage management plan, or engaging in sequential, case-based reasoning.
The present study is positioned to address these limitations directly. By constructing a novel dataset meticulously curated from the foundational psychiatric textbooks described above, we ensure clinical validity and accuracy. By designing eleven distinct and complex QA task types ranging from diagnostic reasoning to treatment planning and sequential analysis, we create a framework to evaluate the sophisticated cognitive functions required in psychiatry. This work, therefore, moves beyond testing what LLMs can do on simplified tasks and begins to assess their readiness for the complex, rigorous, high-stakes reality of clinical mental healthcare.
2.5 Benchmarks for Evaluating LLMs in Psychiatry and Mental Health
Recently, two notable benchmarks illustrate different approaches for mental health and psychiatry. From Empathy to Action: Benchmarking LLMs in Mental Health with MentalBench-10 and a Novel Cognitive-Affective Evaluation Approach [26] proposed a framework that emphasizes empathy, affective alignment, and the capacity of LLMs to translate understanding into actionable guidance. MentalBench-10 focuses on cognitive-affective evaluations, combining both synthetic and user-generated dialogues to assess empathy, appropriateness of advice, and general mental health support. While this contributes valuable insights into the emotional and supportive dimensions of LLM performance, it primarily emphasizes conversational empathy and cognitive-affective alignment rather than the depth of psychiatric reasoning required for clinical tasks.
Similarly, PsychBench: A comprehensive and professional benchmark for evaluating the performance of LLM-assisted psychiatric clinical practice [27] introduced a wide-ranging evaluation platform that incorporates tasks from diagnosis and treatment to professional communication. Its evaluation metrics include ICD Primary Diagnosis Accuracy, Top Choice Alignment Score (TCAS), Medication Match Score (MMS), and Recommendation Coverage Rate (RCR). Although PsychBench benefits from real-world, diverse clinical cases and validation by psychiatrists, its relatively small dataset (300 cases) restricts generalization, and it still relies partly on synthetic dialogue construction, with limited grounding in authoritative, expert-validated psychiatric sources.
Finally, PsychiatryBench is curated from authoritative psychiatric textbooks and casebooks, aiming to provide comprehensive psychiatric reasoning and knowledge evaluation. It introduces eleven distinct QA tasks including diagnostic reasoning, treatment planning, sequential case analysis, and longitudinal follow-up covering more than 5,300 expert-annotated items. The benchmark relies on LLM-as-judge similarity scoring but is built on an expert-validated, broad knowledge base. By grounding its dataset exclusively in clinically validated sources (e.g., DSM-5-TR Clinical Cases, Stahl’s Essential Psychopharmacology), PsychiatryBench directly addresses the gaps left by prior efforts. In contrast to MentalBench-10, which emphasizes affective alignment, and PsychBench, which provides broad but synthetic coverage, our work offers a benchmark that is simultaneously clinically rigorous, multi-task, and extensible.
3 Datasets
(a) Overview of the full PsychiatryBench dataset, including manual extraction and filtering steps.
(b) Structure of the selected subset (S-dataset) used for LLM evaluation.
Figure 1: Visual representation of the PsychiatryBench dataset development pipeline and its final evaluation subset.
The data set used in this study was manually curated to address the unique challenges of psychiatric question answering (QA), with a focus on a diverse range of clinically relevant tasks. Sourced from authoritative psychiatry textbooks and expert-validated clinical resources, the dataset comprises natural language questions paired with expert-formulated answers. These questions are grounded in real-world psychiatric scenarios. They are designed to evaluate the reasoning, decision-making, and knowledge application abilities of LLMs in both clinical and educational contexts.
Figure 1 provides a visual summary of the composition across eleven task categories, while Table 1 outlines the quantitative distribution of samples per task type. Together, they offer a comprehensive overview of the dataset structure and task diversity used in model evaluation.
The manual collection process ensured that the dataset maintains high-quality, medically accurate content, aligning with standardized diagnostic frameworks such as the DSM-5 and ICD-10, as well as clinical practice guidelines from organizations like NICE and the American Psychiatric Association (APA).
The dataset is structured around eleven distinct QA task types, each representing a different cognitive or clinical function critical to psychiatric practice:
  1. 1.
    Diagnosis: This subset contains 467 expert-annotated question-answer pairs centered on diagnostic reasoning. Each item includes a clinical vignette and a diagnostic query, requiring inference of the most appropriate psychiatric diagnosis. The dataset represents 123 unique diagnostic questions and 295 distinct clinical histories. Examples are drawn from six authoritative sources, including 100 Cases in Psychiatry, Geriatric Psychiatry, and Case Files Psychiatry.
  2. 2.
    Treatment: This subset comprises 258 expert-annotated clinical question-answer pairs focused on treatment decision-making in psychiatry. Each example includes a detailed case history, a mental state examination, and a treatment-related clinical question, accompanied by a model answer curated from authoritative psychiatry textbooks. The scenarios reflect real-world psychiatric care and cover a wide range of disorders and treatment modalities, including pharmacological management, psychological therapies (e.g., CBT, psychodynamic therapy), and patient-specific treatment considerations. Most cases are sourced from well-established educational resources such as 100 Cases in Psychiatry, Case Files Psychiatry, and Geriatric Psychiatry.
  3. 3.
    Treatment Follow-Up: This dataset comprises 27 clinical question-answer pairs focused on the follow-up phase of psychiatric treatment. Each scenario includes a psychiatric case history and a primary follow-up question, with a subset (6 entries) extending to a second-tier follow-up. These items evaluate a model’s capacity to support longitudinal clinical care such as medication adjustments, relapse monitoring, and continued patient engagement. All entries were curated from expert-reviewed psychiatric cases to reflect authentic post-treatment scenarios.
  4. 4.
    Classification: This component includes 418 psychiatric case descriptions, each annotated with both a broad diagnostic category and a specific disorder label. The classification task requires models to interpret narrative clinical data and assign appropriate labels by frameworks such as the DSM-5. Cases span multiple psychiatric disorders and are sourced from various clinical and educational texts. This task is crucial for evaluating a model’s ability to operationalize diagnostic taxonomy and differentiate between related disorders.
  5. 5.
    Management Plan: This dataset consists of 337 case-based question-answer pairs that test a model’s ability to generate psychiatric management plans. Each scenario includes a clinical context, often supplemented with mental and physical examination findings. The goal is to evaluate how models handle care coordination, safety planning, treatment adherence, and multidisciplinary strategies. Questions are sourced from authoritative psychiatric references and reflect decisions clinicians make beyond diagnosis or single-treatment choices.
  6. 6.
    Clinical Approach: This subset contains 56 expertly crafted scenarios aimed at evaluating a model’s ability to reason through complex psychiatric cases. The questions focus on the rationale behind clinical decisions such as ruling out differential diagnoses, identifying red flags, or determining the most appropriate investigative or management step. Each case includes multiple LLM-generated responses, enabling comparative analysis. The dataset emphasizes the "how" of clinical reasoning, offering a rigorous benchmark for evaluating the interpretive depth of psychiatric decision support systems.
  7. 7.
    Mental QA: This dataset includes 326 expert-annotated question-answer pairs that test psychiatric knowledge, definition, and conceptual understanding. Most questions 231 explicitly begin with "define," while others explore variations such as "what is," "differentiate," and "describe." Questions are sourced primarily from the Casefile and Case Vignettes series, and include items such as "How would you define thought withdrawal?" or "What are the key diagnostic clues that suggest this patient has histrionic personality disorder?" This dataset targets factual psychiatric knowledge and terminology aligned with DSM-5, ICD-10, and formal psychiatry training curricula.
  8. 8.
    Sequential Question Answering: This dataset comprises 32 clinical vignettes, each followed by a structured sequence of five diagnostic and management questions. The questions reflect a progressive clinical reasoning path, covering: (1) differential diagnosis, (2) supporting evidence and additional needed information, (3) etiological factors, (4) treatment options, and (5) prognosis. The dataset supports multi-turn reasoning and longitudinal clinical understanding. Each case is also tagged with a classification label (32 unique categories), including disorders like schizophrenia, psychotic depression, and substance-related disorders. This structure is ideal for evaluating retention, reasoning, and contextual depth in LLMs.
  9. 9.
    Multiple Choice Questions (MCQs):  This task group encompasses structured assessment formats commonly used in psychiatric exams. It is subdivided into four types of multiple-choice datasets:
    • MCQ:
      1473 standard multiple-choice questions, each with 4-5 answer options. These are clinical knowledge checks grounded in patient histories and symptom-based reasoning. For instance, one item describes a 40-year-old male with choreiform movements and a family history of psychiatric illness, requiring identification of the gene location associated with Huntington’s disease.
    • Exams:
      A curated exam-style dataset, designed to simulate standardized testing disorders. This set is split into two parts for evaluation: EMI 300 and MCQ 300. These items assess factual recall and decision-making across psychiatric domains.
    • EMI (Extended Matching Items):
      A clinical reasoning task where a thematic list of 8-15 options is presented, followed by multiple clinical vignettes requiring matched answers. The EMI format encourages fine-grained differentiation across closely related disorders or treatments. The dataset contains 277 whole EMI clusters, each with shared headers and grouped questions.
    • EMI Separated:
      To support fine-tuning and task-specific evaluation, the full EMI format was disaggregated into 1037 standalone question items. Each retains its original header and option list for context, but functions as an independent unit. This format simplifies scoring and improves input granularity during training.
      Although the original data includes answer explanations, these were excluded in modeling tasks to maintain fairness and prevent leakage.
Each task type is intended to test different facets of psychiatric competence, ranging from factual recall and concept definitions to complex clinical reasoning and longitudinal patient care. This diversity in task design allows for a comprehensive evaluation of both general-purpose and domain-specialized LLMs.
Table 1 presents a matrix of the dataset. Each row corresponds to a specific educational resource (book), and each column represents one of the eleven QA tasks. A checkmark (✓) indicates that the book contains content suitable for that task. For example, Case Files Psychiatry supports nearly all tasks, making it ideal for multi-task learning models. Conversely, MCQs in Psychology focus solely on multiple-choice questions, and the Casebook in Abnormal Psychology is annotated for its relevance to follow-up care and classification.
This matrix provides a clear taxonomy of task-resource alignment, allowing for filtered dataset construction. Researchers can, for example, extract subsets specific to diagnostic reasoning or build MCQ-only datasets for exam simulation. The manual annotation process guarantees transparency, interpretability, and adaptability to various downstream applications.
The curated nature of this dataset, grounded in expert-reviewed content, ensures its relevance to real-world psychiatric practice. It offers a robust and modular benchmark for evaluating LLMs in clinical decision support, medical training, and knowledge-grounded natural language processing. By incorporating diverse task types and detailed evaluation protocols, this dataset provides a comprehensive framework for measuring model performance and guiding improvements in mental health AI.
Table 1: The table shows which psychiatry textbooks and case-based resources include content related to specific clinical tasks. A check mark (✅) indicates coverage of a task, helping identify which resources contribute to each area in the dataset collection.
Books Diagnoses Treatment Treatment Follow-Up Classification Management Plan Clinical Approach Mental QA Sequential QA MCQs
Geriatric psychiatry [28]
Case Files Psychiatry [29]
Clinical Psychology Casebook [30]
DSM-5 Clinical Cases [31]
DSM-5-TR® Self-Exam Questions[32]
100 Cases in Psychiatry [33]
Core Clinical Cases [34]
Casebook in Abnormal Psychology [35]
Revision MCQ and EMI [36]
Psychiatry case vignettes [37]
Clinical Cases in Psychiatry [38]
Stahl’s Psychopharmacology Cases [39]
Stahl’s Essential Cases [40]
MCQs in Psychology [41]
Table 2: Distribution of Available Datasets Categorized by Clinical Task Type
Clinical Task Number of samples
Diagnoses 467
Treatment 258
Treatment Follow-Up 27
Classification 418
Management Plan 337
Clinical Approach 56
Mental QA 326
Sequential Question Answering 32
Multiple Choice Questions 1473
Extended Matching Items 277
Extended Matching Items Separated 1037
Exams 600
Total 5308
4 Experimental Setup
This section outlines the experimental framework used to evaluate the performance of LLMs across a diverse set of psychiatric question-answering (QA) tasks. We describe the eleven task types designed to test specific aspects of psychiatric reasoning and detail the models employed, including both general-purpose LLMs and domain-specialized medical models. The evaluation methodology includes both automatic metrics and LLM-as-a-judge assessments to capture performance on accuracy, reasoning depth, and clinical alignment. By structuring our experiments around realistic clinical tasks and rigorous evaluation criteria, we aim to produce insights into the capabilities and limitations of current LLMs in mental health domains.
4.1 Task Types
The task addressed in this work is a Question Answering (QA) problem in the medical domain, with a specific focus on psychiatry. The dataset consists of natural language questions and expert-curated answers derived from authoritative psychiatry textbooks. These questions span a variety of clinically relevant scenarios and concepts, reflecting the types of reasoning and decision-making tasks encountered in psychiatric practice and education.
Unlike traditional QA systems focused on a single type of question, this dataset encompasses eleven distinct QA tasks, each representing a different cognitive or clinical function in psychiatry. These tasks range from diagnostic reasoning to treatment planning, classification, and sequential case-based analysis. The goal is to evaluate the ability of QA models to understand, reason, and generate medically accurate responses across these varied task types.
Given the complexity and depth of psychiatric knowledge, this work aligns with applications in clinical decision support, medical education, and knowledge-grounded language modeling. The QA challenges posed by this dataset require both factual recall and higher-order reasoning consistent with medical standards such as the DSM-5 and ICD-10.
Below, we briefly describe each of the eleven task types included in the dataset:
4.1.1 Task 1: Diagnoses
This task focuses on determining psychiatric diagnoses based on rich clinical vignettes that simulate real-world case presentations. Each entry in this section is a natural language question paragraph with a detailed patient scenario, often incorporating elements such as the patient’s psychiatric and medical history, presenting symptoms, and sometimes findings from the Mental State Examination (MSE) or contextual physical observations. The cases require models to identify either a likely diagnosis or generate a differential diagnosis, drawing on a nuanced understanding of symptom patterns, temporal progression, and psychosocial factors. For instance, cases may involve individuals with compulsive behaviors and intrusive thoughts indicative of obsessive-compulsive disorder or patients presenting with fear and avoidance consistent with specific phobias. The task challenges models to interpret and synthesize complex narratives, apply diagnostic reasoning, and align answers with formal criteria from standard psychiatric classifications such as the DSM-5 or ICD-10. As such, it plays a critical role in evaluating systems intended for clinical decision support, psychiatric triage, and medical education.
4.1.2 Task 2:Treatment
This task centers on identifying appropriate treatment strategies for various psychiatric disorders based on the clinical scenarios provided. Each case includes a natural language question focused on management, along with contextual patient information such as symptoms, diagnosis, comorbidities, or previous treatment responses. The goal is to generate or select medically accurate treatment plans that align with current psychiatric guidelines. Treatment recommendations may involve pharmacological interventions such as prescribing antidepressants, antipsychotics, mood stabilizers, or anxiolytics, as well as non-pharmacological approaches, including cognitive behavioral therapy (CBT), psychoeducation, or hospitalization in severe cases. In some instances, the model must also account for patient-specific factors, such as age, substance use, risk level, or treatment resistance. This task tests the system’s ability to apply evidence-based psychiatry, reflect clinical guidelines (e.g., NICE, APA), and personalize care based on the patient’s profile. It is particularly relevant for clinical decision support systems, therapeutic recommendation engines, and educational tools aimed at medical trainees.
4.1.3 Task 3: Treatment Follow-Up
This task addresses the longitudinal management of psychiatric patients, focused on follow-up care after an initial treatment decision has been made. This subset of clinical vignettes spans multiple time points, allowing models to reason about how a patient’s mental or physical disorders evolve in response to therapy. Each entry includes an initial clinical scenario, one or more follow-up reports (e.g., new symptoms, lab results, patient response), and a question prompting clinical judgment about how care should be adjusted. These follow-up scenarios often require interpreting medication effects, identifying emerging side effects (e.g., serotonin syndrome, orthostatic hypotension), assessing treatment adherence, or recommending new interventions. In addition to psychiatric symptoms, the task incorporates age-specific considerations (e.g., geriatric psychiatry) and physical health interactions, reflecting the complexity of real-world cases. Models are expected to make informed decisions about whether to continue, adjust, or discontinue treatment. This task simulates ongoing psychiatric care, testing a model’s ability to adapt recommendations over time and handle incomplete, evolving information, making it critical for clinical decision support, chronic care planning, and training in longitudinal psychiatric case management.
4.1.4 Task 4:Classification
A multi-label dataset for mental health disorder classification is created by manually extracting case studies and their corresponding final classifications from multiple books. Each extraction undergoes a review process to ensure quality. Disorder names across different books often show inconsistencies, with similar disorders labeled under varying names. To standardize the terminology, two disorder lists are developed: a category disorders list and a specific disorders list. Both lists are extracted from the DSM-5-TR and ICD-11 manuals and are reviewed by a licensed psychiatrist to ensure accuracy. Disorder names from the books are then mapped to the standardized lists using Gemini 2.5 Pro. This mapping process is conducted twice, once using the disorder categories list and once using the specific disorders list. As a result, each case in the final dataset is assigned two sets of disorder labels: one set corresponding to the specific disorders and another set corresponding to their broader categories.
4.1.5 Task 5: Management Plan
This task involves generating a structured and contextually appropriate management plan for patients presenting with psychiatric symptoms. Each entry in this portion’s detailed clinical history is often accompanied by information from a Mental State Examination (MSE) and Physical Examination, followed by a question such as "How would you manage this case?" or "What interventions should you offer?". The answers outline holistic, step-by-step care strategies, integrating diagnostic clarification, risk assessment, therapeutic planning, and follow-up. Management plans may include pharmacological treatments (e.g., SSRIs for generalized anxiety), psychological therapies (e.g., CBT for OCD), psycho-education, social interventions, and referral pathways. The task requires models to interpret multiple layers of clinical input and propose management approaches aligned with psychiatric guidelines (e.g., NICE or APA). It emphasizes the practical application of knowledge in real-world clinical decision-making and supports training for complex care coordination in psychiatric settings.
4.1.6 Task 6: Clinical Approach
This task focuses on evaluating the appropriate clinical approach to a variety of psychiatric presentations, emphasizing step-by-step assessment and decision-making processes. Each question is tied to a case scenario and asks how clinicians should approach diagnosis, risk assessment, or management in a structured and evidence-based manner. Unlike tasks that ask for final diagnoses or treatments, this task emphasizes process-oriented reasoning on how clinicians gather information, rule out critical disorders, prioritize differential diagnoses, and choose appropriate initial interventions. The dataset features cases such as evaluating major depressive disorder, schizophrenia, panic disorder, or suicidal, with questions often framed as "What is the clinical approach to…?". Answers are grounded in established psychiatric practice guidelines and typically reference core elements such as history taking, Mental State Examination (MSE), collateral information, risk stratification, and initial management priorities. This task simulates the real-world cognitive workflow of psychiatric professionals, making it essential for models designed to assist with clinical decision-making, medical education, and case-based simulation training in psychiatry.
4.1.7 Task 7: Mental QA
This task involves the definition and clarification of core psychiatric concepts, syndromes, and clinical terms. Each question in the dataset prompts a concise and medically accurate explanation of a term or concept relevant to mental health, such as "chelonian," "bizarre delusions," or the pharmacological class "selective serotonin reuptake inhibitors." These terms are foundational to understanding clinical psychiatry and are often encountered in diagnostic interviews, treatment planning, and academic contexts. The answers are short, definitional statements drawn from clinical knowledge and psychiatric literature, consistent with sources like the DSM-5, pharmacological references, or foundational psychiatry textbooks. This task is designed to enhance conceptual grounding for QA models and promote explainability, particularly in systems intended for clinical education, patient-facing applications, or knowledge base construction. It also supports retrieval-augmented models and interpretable AI by encouraging explicit reasoning about medical terminology.
4.1.8 Task 8: Sequential Question Answering
This task simulates dynamic, case-based psychiatric reasoning through a structured sequence of interrelated questions. Each scenario presents a detailed patient history, followed by a series of clinical queries that must be answered in context. The questions are designed to mirror the natural flow of psychiatric evaluation, moving from the initial differential diagnosis to evidence-based justification, etiological analysis, treatment planning, and finally prognostic estimation. For instance, a single case might begin by asking "What is the likely differential diagnosis?" followed by questions like "What information supports the diagnosis?", "What are the important etiological factors?", and "What treatment options are available?". This multi-turn format requires the model to maintain coherence, track evolving information, and apply lay clinical reasoning over time. It is beneficial for developing and evaluating QA systems that simulate real-world psychiatric interviews, medical board exam preparation, or decision-support tools in longitudinal patient care. This task emphasizes temporal context, cognitive continuity, and adaptive inference, making it a high-value benchmark for sequential understanding in medical AI.
4.1.9 Task 9: Multiple Choice Questions (MCQ)
This task includes three widely used assessment formats in psychiatric education and certification: (1) standard multiple-choice questions (MCQs), (2) extended matching items (EMIs), and (3) full-length exam simulations. Together, they are designed to evaluate knowledge recall, clinical reasoning, and decision-making in structured, test-like disorders.
  1. 1.
    Multiple Choice Questions (MCQ): The standard MCQs in our dataset are drawn from psychiatry-focused exams and cover a broad range of topics, including diagnosis, treatment, symptom classification, theoretical models, and psychopharmacology. Each question presents a clinical or conceptual prompt, followed by a list of five options, with one correct answer. These questions are designed to assess factual knowledge as well as applied understanding and clinical judgment.
  2. 2.
    Extended Matching Items (EMI): The EMI format groups several related questions under a single theme, such as diagnostic systems or therapeutic categories. Each item shares a standard answer list and must be matched with the most appropriate response. To maximize the utility of this format for QA evaluation, we split each EMI set into separate question-answer pairs while maintaining the original answer pool context. Enables fine-grained evaluation of the model’s ability to differentiate between similar options and interpret questions that depend on thematic nuance. The EMI task is particularly effective at testing classification accuracy, thematic grouping, and structured reasoning across semantically related content.
  3. 3.
    Exam Simulations: We also included complete exam sets composed of multiple MCQs delivered together, replicating the structure and pacing of real psychiatric board exams. These exams are designed to simulate actual test disorders, including variation in difficulty, topic coverage, and question style. This component allows for evaluation of performance across a complete test and supports use cases such as automated exam preparation and curriculum-aligned feedback systems.
Taken together, these three formats provide a comprehensive benchmark for evaluating model performance on structured psychiatric assessments. They also support practical applications in educational settings, such as intelligent tutoring, exam readiness, and adaptive learning.
4.2 Language Models
In this section, we detail the LLMs evaluated throughout our study, covering both general-purpose and domain-specialized systems. The goal is to assess their capabilities in handling various clinically relevant tasks, such as diagnosis, treatment planning, and psychiatric reasoning. By comparing general LLMs with medical-domain-specific models, we aim to highlight the performance trade-offs and advantages of each category. This dual perspective provides a clearer understanding of how LLMs can be optimized or selected for different healthcare applications.
4.2.1 General Models
Table 3: Summary of LLMs used in our experiments
Model Size (Parameters) Source API Service Provider
Gemini 2.5 Pro Preview 03-25 [42] Proprietary Closed-source Google’s Gemini API
Gemini 2.5 Flash Preview 04-17 (Thinking) [42] Proprietary Closed-source Google’s Gemini API
Gemini 2.5 Flash Preview 04-17 [42] Proprietary Closed-source Google’s Gemini API
Gemini 2 Flash [42] Proprietary Closed-source Google’s Gemini API
Claude Sonnet 4 [43] Proprietary Closed-source Anthropic API
DeepSeek R1 [44] 236B Open-source DeepSeek API
DeepSeek Chat [44] 67B Open-source DeepSeek API
LLaMA 3.3 70B [45] 70B Open-source Together API
QWQ 32 [46] 32B Open-source Nvidia API
QWEN 3 32b [46] 32B Open-source Deepinfra API
Table  3 presents a concise overview of the LLMs evaluated in the study. It categorizes each model by name, parameter size, source type (open or closed), and the API service provider used to access them. The models include both proprietary and open-source systems. Proprietary models, such as various versions of Google’s Gemini (e.g., Gemini 2.5 Pro, Flash Preview, and Flash), and Anthropic’s Claude Sonnet 4, are closed-source, meaning their architectures and training details are not publicly available. These models are accessed through their respective API services, such as Google’s Gemini API and the Anthropic API.
On the other hand, open-source models like DeepSeek R1 (236B), DeepSeek Chat (67B), LLaMA 3.3 70B, QWQ 32, and QWEN 3 32B offer more transparency in terms of model architecture and training datasets. These are made available through platforms such as the DeepSeek API, Together API, Nvidia API, and Deepinfra API.
The parameter sizes of the open-source models are explicitly stated, which is important for understanding their potential capabilities and computational requirements. The proprietary models do not disclose their parameter counts, reflecting the limited public insight into their underlying architectures.
Overall, the table illustrates a diverse selection of LLMs sourced from both major tech companies and open research communities, providing a broad basis for comparative experimentation and analysis.
4.2.2 Medical Models
Table 4: Comparison of Medical Language Models
Model Size (Parameters) Source API Service Provider
Llama3-OpenBioLLM-70B [47] 70B Open-source Hugging Face
Palmyra-Med-70B [48] 70B Open-source Hugging Face
Llama-MedX v3.2 [49] 70B Open-source Hugging Face
JSL-MedLlama-3-8B-v2.0 [50] 8B Open-source Hugging Face
MedGemma [51] 27B Open-source Hugging Face
II-Medical-8B [52] 8B Open-source Hugging Face
Med-PaLM (v2) [53] Proprietary Closed-source Google Cloud Vertex AI
To ensure a well-rounded and insightful evaluation of LLMs in the context of psychiatric question answering, we selected a diverse group of medical-focused models that reflect a range of architectures, parameter scales, and accessibility levels. This curation was designed to balance frontier performance with practical considerations such as openness, reproducibility, and ease of deployment.
Our model set includes high-capacity open-source systems like Llama3-OpenBioLLM-70B, Palmyra-Med-70B, and Llama-MedX v3.2, which leverage recent advances in foundational models and are optimized for biomedical and clinical tasks. We also included JSL-MedLlama-3-8B-v2.0, a smaller model aimed at providing strong performance while remaining lightweight and adaptable. These open-source models are readily accessible via platforms such as Hugging Face and are suitable for both academic research and fine-tuned clinical deployments.
In contrast, we also evaluated Med-PaLM (v2), a proprietary model developed by Google DeepMind. While not open-source, Med-PaLM represents a high-performing benchmark for medical reasoning and is accessible through Google Cloud’s Vertex AI platform. Its inclusion provides an important comparison point for understanding the trade-offs between commercial-grade deployment and the flexibility of open systems.
By including both open and closed models of varying sizes, we aim to assess not only raw accuracy on psychiatric tasks but also broader factors such as infrastructure requirements, cost, adaptability, and transparency, key considerations for real-world clinical and educational use. A summary of these models, along with their sources and access points, is provided in Table 4.
4.3 Evaluation Methodology
To assess the performance and alignment of model outputs, we employ a set of evaluation metrics designed to quantify agreement, consistency, and quality across different prompts and models. These metrics allow us to systematically compare outputs, identify patterns in model behavior, and measure improvements across prompt variations. In particular, we focus on agreement counts between model pairs to capture alignment and shared reasoning under different prompting disorders.
4.3.1 Accuracy Metrics
We mainly use two accuracy metrics in this work;
  • Balanced Accuracy (BA) is used to handle imbalanced datasets by computing the average recall for each class. It is defined as follows:
    $BA=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)$ (1)
    BA helps in evaluating the model’s performance when the class distribution is imbalanced.
  • Accuracy is the proportion of correctly classified samples among the total number of samples:
    $Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (2)
    Accuracy gives an overall measure of correctness, but it may not be reliable in cases of class imbalance, as it can be biased towards the majority class.
Here, $TP$ (True Positives) and $FN$ (False Negatives) are used to compute recall (sensitivity), while $TN$ (True Negatives) and $FP$ (False Positives) are used to compute specificity.
4.3.2 Extended Matching Item Evaluation
In the EMI set, each question consists of multiple subquestions, all of which must be answered based on shared response options. To ensure accurate data evaluation, the dataset was assessed in two distinct formats: first, as a complete EMI set (full-format evaluation), and second, by treating each subquestion as an independent query with its evaluation (separated format).
In the full EMI set format, scoring follows the Partial Scoring System (PCS), which assigns partial credit based on the proportion of correctly answered subquestions in a set. In contrast, the separated subquestion format evaluates each subquestion independently using a binary scoring system.
For both formats, any instance of multiple answers being given for a single question is treated as incorrect and scored as zero.
1. Partial Scoring System (PCS) - Full EMI Set Evaluation
This evaluation method measures how well the model performs on grouped EMI questions, where each set consists of multiple related subquestions. The scoring system allows for partial credit, ensuring that models are not penalized excessively for a few incorrect subanswers.
The score for the $i$-th EMI set is computed as the proportion of correct subquestions:
$EMI\_Score_{i}=\frac{1}{S_{i}}\sum_{j=1}^{S_{i}}c_{ij}$ (3)
where $N$ is the total number of EMI sets, $S_{i}$ is Number of subquestions in the $i$-th EMI set, and $c_{ij}\in\{0,1\}$: Binary indicator for whether the $j$-th subquestion in the $i$-th set is answered correctly.
The overall EMI accuracy is the average score across all sets:
$EMI\_Accuracy=\frac{1}{N}\sum_{i=1}^{N}EMI\_Score_{i}$ (4)
2. Individual Subquestion Evaluation - Separated EMI
This approach evaluates each subquestion independently, allowing for standard binary scoring per subquestion. It offers a fine-grained view of model performance without being affected by the structure of the EMI sets.
Let:
  • $T$: Total number of individual subquestions across all EMI sets,
  • $a_{k}\in\{0,1\}$: Binary indicator of correctness for the $k$-th subquestion when evaluated independently.
The overall accuracy is then computed as:
$\text{Subquestion\_Accuracy}=\frac{1}{T}\sum_{k=1}^{T}a_{k}$ (5)
Answer Integrity Rule - Penalty for Multiple Answers
In both evaluation methods, if the model produces multiple answers for a single question (whether set-based or individual), the answer is considered invalid and marked as incorrect. This ensures clarity, fairness, and aligns with human assessment standards.
$c_{ij}=0\quad\text{if multiple answers are provided}$ (6)
Common Correct: This measures the number of questions answered correctly in both formats:
$\text{CommonCorrect}=\sum_{i=1}^{N}\mathbf{1}\left(F_{i}=1\land S_{i}=1\right)$ (7)
where $F_{i}$ and $S_{i}$ are indicator variables for Full format response and Separated format response on question $i$, respectively. $F=1$ if the model answers question $i$ correctly in the Full format and 0 otherwise. $S_{i}=1$ if the model answers question $i$ correctly in the Separated format and 0 otherwise. $\land$: The logical AND operator.
Common Incorrect: This measures the number of questions answered incorrectly in both formats:
$\text{CommonIncorrect}=\sum_{i=1}^{N}\mathbf{1}\left(F_{i}=0\land S_{i}=0\right)$ (8)
Overall Consistency: This score reflects how often the model yields the same result (correct or incorrect) across both formats:
$\text{Consistency}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\left(F_{i}=S_{i}\right)$ (9)
Format Agreement Rate: This metric calculates the proportion of questions for which the model was correct in both formats, relative to all cases where the model was correct in at least one:
$\text{Agreement}=\frac{C}{C+O_{F}+O_{S}}$ (10)
where $C$ is the number of the correct questions in both formats, $O_{F}$ are the correct questions only in Full format, and $O_{S}$ are the correct questions only in the separated format.
Format Divergence Rate: This captures how often the model’s correctness depends on format. A high divergence suggests inconsistent reasoning between question structures:
$Divergence=\frac{O_{F}+O_{S}}{N}$ (11)
These consistency metrics are crucial for evaluating the robustness and fairness of models, especially in educational and clinical settings where presentation format should not influence correctness.
4.3.3 LLM as a judge
To quantitatively evaluate the output quality of model-generated answers against ground truth references, we employed an automated approach leveraging an LLM as an expert judge [54]. Specifically, Llama 3 70B was configured for this evaluation task. The core objective was to compute a similarity score, ranging from 0 to 100, representing the degree of substantive congruence between the model-generated response (referred to as the Candidate Answer) and the reference ground truth (referred to as the Reference Answer). Importantly, the evaluation was designed to focus on the core clinical content, including reasoning, correctness, and completeness of the response while disregarding differences in writing style or phrasing. 10.3
Recognizing the sensitivity of LLM evaluations to prompt design, we experimented with two distinct prompts, generated by tasking two capable LLMs, specifically GPT-4.5 and gemini-2.0-pro-exp-02-05, to create instructions for calculating the design similarity score (0–100). The prompt generation strategies differ: for GPT-4.5, we guided it toward a score-based prompt, employing a point-based system for different content components. Conversely, for gemini-2.0-pro-exp-02-05, we aimed for a more logic-driven prompt that relied on illustrative examples (few-shot learning) and emphasized understanding the overall structure and meaning of the Reference and Candidate Answers.
The prompt generated by GPT-4.5 is shown in Figure 15, and the prompt generated by gemini-2.0-pro-exp-02-05 is shown in Figure 16. After comparative testing of both prompts with the Llama 3 70B judge, we selected the gemini-2.0-pro-exp-02-05-generated prompt for our final evaluation pipeline. This decision was driven by the observation that the logic-driven, example-based prompt elicited more consistent and reliable similarity scores across a diverse range of clinical tasks and model responses. These tasks included diagnostic assessment, treatment planning, follow-up care, management strategies, clinical approaches, mental health question-answering, and multi-turn clinical reasoning. The example-based prompt aligned more closely with our focus on nuanced content evaluation, as compared to the more rigid point-based system used in the GPT-4.5 prompt.
Once individual similarity scores $S_{i}\in[0,100]$ are obtained for each task instance $i$, we compute the final evaluation score by taking the arithmetic mean across all evaluated instances:
$\text{Final Score}=\frac{1}{N}\sum_{i=1}^{N}S_{i}$ (12)
where $N$ is the number of datasets used. This final score serves as a robust metric for quantifying the overall content-level alignment between model-generated outputs and expert-authored references across multiple clinical domains.
4.3.4 Classification Evaluation
Since the classification task is multi-label in nature, each instance may be associated with multiple correct labels. To accurately assess model performance in this setting, we employed two complementary evaluation metrics: Subset Accuracy and Weighted-Average F1-Score. These metrics together offer a balanced view rewarding exact correctness while also accounting for partial matches and class imbalance. [55]
Subset Accuracy (Exact Match Ratio)
Subset Accuracy is a stringent metric that only considers a prediction correct if the predicted label set exactly matches the true label set no missing or extra labels are allowed. This makes it particularly valuable when precision in full-label prediction is critical. It answers the question: "How often did the model get everything exactly right?"
$\text{Subset Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{I}(\hat{Y}_{i}=Y_{i})$ (13)
Where: $N$ is the total number of samples, $Y_{i}$ is the set of true labels for the $i$-th sample, $\hat{Y}_{i}$ is the set of predicted labels, $\mathbf{I}(\cdot)$ is the indicator function, and equal to 1 if $\hat{Y}_{i}=Y_{i}$, and 0 otherwise.
Weighted-Average F1-Score (F1-weighted)
While Subset Accuracy requires perfect prediction, the F1-weighted score provides a more flexible and balanced view of model performance. It evaluates how well the model balances precision (avoiding false positives) and recall (avoiding false negatives) for each label independently. Then, it aggregates these per-label F1-scores into a single number using a weighted average, where more common labels contribute more to the final score.
First, the F1-score for each label $l$ is calculated as the harmonic mean of precision ($P_{l}$) and recall ($R_{l}$):
$F_{1,l}=2\times\frac{P_{l}\times R_{l}}{P_{l}+R_{l}}$ (14)
Then, the overall weighted-average F1-score across all labels $\mathcal{L}$ is computed as:
$\text{F1-weighted}=\sum_{l\in\mathcal{L}}w_{l}\cdot F_{1,l}$ (15)
Where $\mathcal{L}$ is the set of all labels, and $w_{l}$ is the proportion of samples in which label $l$ appears in the ground truth.
This metric ensures that performance on frequent labels contributes more heavily, helping to mitigate distortions caused by class imbalance.
4.4 Parsing Model Outputs
The process of interpreting and extracting meaningful information from LLM outputs emerged as a critical and non-trivial component of our methodology. While it may appear straightforward for tasks involving simple responses such as multiple-choice questions (MCQs) or numerical ratings the actual output behavior of large models introduced substantial complexity.
Advanced reasoning-oriented models, such as DeepSeek Chat and Gemini Thinking, often diverged from strict output formats. Rather than outputting a clean label or value, they frequently embedded the answer within lengthy justifications or explanations. This behavior required parsing not just the final answer, but also understanding the structure of the response and retaining the embedded reasoning what we refer to as "printing the reasoning".
Consequently, our methodology included customized parsers and heuristics tailored to each model’s output style, allowing us to isolate the relevant information while preserving contextual evidence to support qualitative evaluation.
5 Methodology Implementation
This section outlines the comprehensive methodology used to develop, refine, and evaluate the PsychiatryBench dataset a benchmark designed to assess the clinical reasoning capabilities of LLMs in psychiatry. We describe the end-to-end pipeline, beginning with the manual curation of high-quality psychiatric question-answer (QA) pairs from evidence-based sources, followed by rigorous filtering and formatting procedures to ensure clinical accuracy and educational value. We then present the framework used to evaluate model responses across a diverse set of task types, including both structured and open-ended formats. Finally, we detail the design and implementation of specialized tasks that test distinct dimensions of psychiatric competence, such as classification, clinical reasoning, and treatment planning. This modular and scalable methodology provides a robust foundation for benchmarking LLM performance in real-world psychiatric assessment scenarios.
5.1 PsychiatryBench Development Pipeline
Figure 2: Workflow for the PsychiatryBench study: manual extraction and processing of clinical QA samples followed by LLM evaluation across eleven task types.
Our methodology begins with systematic manual extraction of content from standard evidence-based psychiatry textbooks. Subject matter experts thoroughly review and select clinical material with high educational and diagnostic value, vibrant case studies, descriptive patient scenarios, and disorder-specific treatment information. From this material, a diverse set of question-answer (QA) pairs is constructed to represent real-world psychiatric reasoning. The QA items are crafted across eleven core task types: Diagnosis, Treatment, Mental QA, Sequential QA, Management Plan, Treatment Follow-up Clinical Approach, Classification, Multiple Choice Questions (MCQ), and Extended Matching Items (EMI). These task types were chosen for their ability to reflect lay clinical reasoning rather than rote knowledge recall.
Following construction, the QA items are passed through a rigorous filtering process designed to ensure clarity, clinical validity, and educational utility. Ambiguous, overly simple, or redundant items are removed, and the remaining questions are reviewed for content balance and diagnostic diversity. These filtered QA items form the foundation of the PsychiatryBench dataset. In this phase, questions are sorted by task type and formatted consistently, allowing for structured downstream evaluation. A comprehensive data cleaning step is performed next, standardizing answer styles, resolving typographic issues, and ensuring uniform metadata. The result is a refined, high-quality QA dataset that tests not only factual recall but also diagnostic reasoning, treatment planning, and sequential clinical decision-making in psychiatry.
To evaluate the performance of LLMs, we apply each model to the finalized PsychiatryBench dataset. Each model is prompted to generate answers across all eleven selected task types. Importantly, in addition to serving as QA generators, LLMs are also used as evaluators or judges, scoring the generated responses based on accuracy, completeness, and clinical relevance. This dual use of LLMs helps automate and scale the evaluation process while maintaining consistency across judgments. A structured rubric is applied to produce a final score for each model, capturing performance across various clinical dimensions. This evaluation framework enables a robust and scalable benchmark for assessing how well LLMs understand, reason through, and respond to complex psychiatric questions, grounded in authentic medical education content.
5.2 Filtering, Refinement, and Data Cleaning
Following construction, the QA items are passed through a rigorous filtering process to ensure clarity, clinical validity, and educational utility. Ambiguous, overly simplistic, or redundant items are systematically removed. The remaining questions are reviewed to ensure a balance of content and a diversity of represented diagnoses. These filtered QA items form the foundation of the PsychiatryBench dataset. In this phase, questions are sorted by task type and formatted consistently to allow for structured downstream evaluation. A comprehensive data cleaning step is then performed, which involves standardizing answer styles, resolving typographic issues, and ensuring all metadata is uniform. The result is a refined, high-quality QA dataset that tests not only factual recall but also diagnostic reasoning, treatment planning, and sequential clinical decision-making in psychiatry.
5.3 Prompting Strategy
All models were evaluated using a standardized zero-shot prompting approach. While the same prompt format was applied uniformly across all models to ensure fairness in comparison, prompt design was tailored to each specific task type within PsychiatryBench. For example, diagnostic reasoning tasks employed structured clinical case prompts, whereas factual knowledge tasks like Mental QA used direct definition queries. This task-specific prompting allowed us to preserve the integrity and intent of each evaluation domain while maintaining cross-model consistency.
5.4 LLM Models Implementation
We implemented and evaluated a total of 11 LLMs, comprising both general-purpose frontier models (e.g., Gemini 2.5 Pro, Claude Sonnet 4, LLaMA 3.3, QWQ-32) and medical domain-specific models (e.g., MedGemma, OpenBioLLM). These models were tested across all eleven PsychiatryBench task types to assess their performance in real-world psychiatric reasoning tasks.
Each model was prompted using standardized engineering and evaluation templates tailored to the specific requirements of each task. Model outputs were evaluated using both automatic metrics (e.g., accuracy, F1 score) and an LLM-as-a-judge framework, which relied on a reference model to score responses based on semantic similarity and clinical relevance. This dual evaluation strategy ensured consistency, accuracy, and clinical alignment across diverse task types.
5.5 LLM Performance Evaluation Framework
To evaluate the performance of LLMs, we apply each model to the finalized PsychiatryBench dataset. Each model is prompted to generate answers across all eleven selected task types. Our evaluation framework employs a multi-faceted approach tailored to the specific nature of each task. For open-ended reasoning tasks such as Diagnosis, Treatment, and Management Plan, we utilize an LLM-as-a-judge methodology. In this dual-role capacity, a designated LLM scores the generated responses from other models based on accuracy, completeness, and clinical relevance, providing a nuanced measure of semantic alignment. For structured tasks, we apply specific quantitative metrics. For instance, the Classification task is evaluated using Subset Accuracy and F1-scores to handle its multi-label nature. At the same time, multiple-choice and Extended Matching Items are scored using accuracy and custom partial scoring systems. This hybrid evaluation framework, combining qualitative LLM-based scoring with task-specific quantitative metrics, enables a robust and scalable benchmark for assessing how well LLMs understand, reason through, and respond to complex psychiatric questions. To ensure consistency, all generated outputs were evaluated using a single evaluator model (LLaMA 3.3 70B) across all tasks. Each output and its reference were scored using a prompt-generated similarity scoring rubric. The evaluator model remained constant throughout to avoid bias introduced by evaluator variation.
5.6 Task-Specific Implementations
To capture the diverse skills required for comprehensive psychiatric reasoning, we developed five specialized task implementations within the PsychiatryBench framework. Each task is designed to probe a distinct facet of clinical expertise from high-level categorization of disorders to stepwise diagnostic reasoning, from rapid knowledge recall in board-style questions to complex pattern recognition in extended matching formats. Below, we outline the goals, data sources, construction workflows, and evaluation criteria for each task type: Classification, Clinical Approach, Multiple Choice Questions (MCQ), and Extended Matching Items (EMI). This modular design enables targeted analysis of model strengths and weaknesses across different dimensions of psychiatric decision-making.
5.6.1 Classification Task
The classification component of the PsychiatryBench was developed to evaluate the ability of LLMs to categorize psychiatric disorders based on clinical descriptions, symptom clusters, and established diagnostic systems such as the DSM-5 and ICD-10. Rather than requiring exact diagnoses, this task involved structured prompts asking models to assign each case to appropriate diagnostic categories (e.g., mood disorders, psychotic disorders) and specific disorders (e.g., bipolar I disorder, schizophrenia). Clinical case samples were manually extracted from authoritative psychiatric casebooks and annotated with preliminary disorder labels. In parallel, two reference lists were created from the DSM-5-TR: one for high-level diagnostic categories and another for specific disorders. A licensed psychiatrist validated both lists to ensure clinical accuracy and alignment with formal diagnostic criteria. A mapping process was then applied to align each sample with both the validated category-level and disorder-level labels. This dual-labeling schema enabled multi-label classification and allowed models to be evaluated on their ability to reason through hierarchical diagnostic structures. Although the classification dataset was ultimately excluded from the final model evaluation to maintain focus on more reasoning-intensive tasks, it played a critical role in shaping the design and validation of other task types, such as Diagnosis, Mental QA, and Management Plan. It also provided key insights into how LLMs interpret psychiatric terminology, differentiate between overlapping disorders, and apply structured diagnostic logic capabilities that are foundational to clinical psychiatric reasoning.
5.6.2 Clinical Approach Task
The Clinical Approach task in the PsychiatryBench is designed to evaluate a model’s ability to emulate the diagnostic thought process of a psychiatrist. Unlike other tasks that target direct question-answering, this component focuses on assessing logical reasoning, hypothesis formulation, and sequential clinical interpretation based on patient case narratives. The process begins with the creation of patient history scenarios, which are either extracted from psychiatric textbooks or crafted to simulate real-world clinical cases. These cases are then used to generate open-ended clinical reasoning questions. Question generation is performed using the Gemini 2.5 Pro model, which is prompted to formulate questions that require interpretive and multi-step reasoning. Each generated question undergoes a manual validation step. Questions that are deemed clinically relevant and educationally meaningful proceed to the next phase; otherwise, they are revised or rewritten to ensure clinical quality. Once validated, the finalized cases and questions are presented to multiple LLMs. Each model is tasked with generating structured, reasoned responses to the open-ended clinical prompts. These responses are not evaluated solely for factual correctness, but also for the quality of clinical logic, appropriateness of diagnostic pathways, and overall coherence of the suggested approach. To explore the effect of prompting methodology, two parallel evaluation tracks were conducted. In the first, questions were generated using the Gemini model. In the second, the same patient cases were used without the question-generation step, allowing for a direct comparison of LLM performance with and without prompt engineering. Each model’s output was scored across criteria of reasoning depth, clinical relevance, and judgment quality, yielding a final score. This task is essential in assessing how well LLMs can replicate the nuanced reasoning process of trained psychiatrists and highlights their potential role in clinical decision support systems.
5.6.3 Multiple Choice Question (MCQ) Task
The Multiple Choice Questions (MCQ) component was developed as part of the broader PsychiatryBench dataset to reflect the structured, high-efficiency testing style commonly used in psychiatric board examinations and educational settings. These questions typically consist of a question stem followed by four to five answer options, with only one correct choice. MCQs are widely used to assess knowledge recall, clinical recognition, and treatment matching, making them a familiar tool for both learners and educators in psychiatry.
The initial pool of MCQs was manually curated from reputable psychiatry resources, including standardized exam preparation materials, question banks, and textbook-based assessments. Many of these original items were accompanied by detailed explanations or rationales justifying why the correct answer was appropriate and why distractors were incorrect. While this information is pedagogically useful, it introduces bias when used in LLM evaluation settings, where models may memorize answer justifications or be influenced by phrasing patterns. To prevent this, a dedicated preprocessing step was implemented to remove all explanatory text from each MCQ. Ensures that only the question stem and multiple-choice options were retained, creating a clean, unbiased format suitable for model interaction and evaluation.
In addition to removing explanations, all answers were standardized. Involved converting varying answer label formats (e.g., Roman numerals, complete sentences, abbreviated codes) into a consistent A/B/C/D labeling scheme. Correct answers were annotated in a separate metadata field to facilitate scoring and model training. Furthermore, ambiguous or duplicate questions were filtered out during this process to maintain quality and reduce redundancy. Although MCQs were excluded from the final model evaluation phase in favor of tasks requiring open-ended reasoning and multi-step logic, they were retained as part of the broader dataset for auxiliary evaluation and pretraining. The MCQ component also served as a structural foundation for the more complex Extended Matching Items (EMI) task, which builds on similar principles but with greater contextual complexity.
5.6.4 Extended Matching Items (EMI) Task
The Extended Matching Items (EMI) component was developed as part of the PsychiatryBench dataset to emulate a standard format used in clinical assessments, particularly within medical and psychiatric education. EMI questions are structured to assess deeper clinical reasoning by presenting a list of possible answer options (the theme), followed by multiple clinical scenarios that require selecting the most appropriate answer from that list.
Each EMI question set typically consists of the following elements:
  • A theme (classes of antidepressants, types of psychotherapies)
  • An option list (usually 8–10 related choices)
  • Several clinical vignettes, each requiring the selection of the most suitable answer from the option list
This format is handy for testing pattern recognition, differential diagnosis, and treatment decision-making in psychiatry. For instance, an EMI set might include different patient cases presenting with depressive symptoms, and the task would be to select the most appropriate treatment or classification subtype for each case. In our methodology, EMI questions were manually created or adapted from authoritative psychiatry textbooks by expert reviewers. Ensures that each item was clinically relevant, unambiguous, and reflected real-world psychiatric decision-making. The EMI format requires the language models not only to recognize correct terminology but also to apply contextual reasoning.
6 Results & Discussion
This section presents the results of our experimental evaluation across multiple psychiatric QA tasks, using a range of LLMs assessed by expert-generated reference answers and LLM-as-a-judge scoring. We analyze performance trends across models and task types, comparing output quality under different prompt strategies and exploring the strengths and weaknesses of each model in tasks such as diagnosis, treatment planning, and follow-up care. Our discussion addresses both quantitative outcomes and qualitative patterns in reasoning, highlighting areas where LLMs demonstrate promising clinical utility and where challenges remain in achieving safe, interpretable, and consistent psychiatric decision support.
6.1 Diagnosis
Table 5: Diagnosis Results: Average Similarity Scores for Model-Generated Diagnosis Outputs.The underlined bold values denote the best score.
Model Name Avg. Similarity score
Deepseek-R1 $84.2$
Deepseek-chat $82.7$
Gemini 2.5 Pro Preview (03-25) 86.0
Gemini 2.0 Flash $83.1$
Llama 3.3 70B $83.6$
QWQ-32 $85.9$
Gemini 2.5 Flash Preview (04-17, Thinking) $85.0$
Gemini 2.5 Flash Preview (04-17) $78.0$
We generated model outputs using the Engineering prompt template 1 and the Evaluation prompt template 2 separately.As shown in Table 6, the average similarity scores, evaluated by a Llama 3.3 70B judge, are reported for eight models on the diagnosis task. The scores range from (78.0%) to (86.0%), indicating that all evaluated models demonstrate strong semantic alignment under the given prompts.The top-performing model, Gemini 2.5 Pro Preview (03-25), achieved the highest similarity score at (86.0%), highlighting the impact of recent architectural refinements or expanded training data. Close behind is QWQ-32, scoring (85.9%), effectively tying for second place and proving competitive despite limited adoption. Gemini 2.5 Flash Preview (04-17, Thinking) followed with a solid (85.0%), showing how variations in inference behavior (e.g., "Thinking" mode) can yield measurable gains in semantic alignment. At the other end of the spectrum, Gemini 2.5 Flash Preview (04-17), without the "Thinking" mode, recorded the lowest score at (78.0%), underperforming relative to its sibling variants. Deepseek-chat scored (82.7%), placing it above the lowest performer but still trailing the top contenders. Mid-tier results included DeepSeek-R1 (84.2%), Llama 3.3 70B (83.6%), and Gemini 2.0 Flash (83.1%), all delivering solid, if not leading, performance. Overall, the results span a compact range of 8.0 percentage points, from (78.0%) to (86.0%), indicating a generally strong level of performance across all models. However, the standout scores from newer or enhanced models, especially Gemini 2.5 Pro Preview, underscore the benefits of continual optimization in both prompt design and model development.
6.2 Treatment
Table 6: Treatment Results: Average Similarity Scores for Model-Generated Treatment Outputs. The underlined bold values are the best score.
Model Name Avg. Similarity score
Deepseek-R1 79.9
Deepseek-chat 81.9
Gemini 2.5 Pro Preview (03-25) 80.5
Gemini 2.0 Flash 77.4
Llama 3.3 70B 82.3
QWQ-32 82.1
Gemini 2.5 Flash Preview (04-17, Thinking) 82.9
Gemini 2.5 Flash Preview (04-17) 81.0
We generated model outputs using the Engineering prompt template 17 and the Evaluation prompt template 18 separately.As show in table  5 all eight models deliver moderately high similarity scores for their treatment recommendations, ranging from (77.4 %) to (82.9 %). The Gemini 2.5 Flash Preview (04-17, Thinking) led the set at (82.9 %), followed closely by Llama 3.3 70B at (82.3 %) and QWQ-32 at (82.1 %). Deepseek-chat achieved (81.91 %), while the secondary Gemini 2.5 Flash Preview (04-17) variant score (81.0 %). Gemini 2.5 Pro Preview (03-25) and Deepseek-R1 recorded (80.5 %) and (79.9%), respectively, with the original Gemini 2.0 Flash trailing at (77.4 %). The spread of (5.49 %) between the top and bottom performers highlights a meaningful performance gap in generating treatment plans. The clear advantage of the Gemini 2.5 Flash Preview (04-17, Thinking) suggests that its latest reasoning capabilities particularly enhance clinical coherence and relevance. Llama 3.3 70B and QWQ-32 also demonstrate that robust language understanding and contextual breadth translate into more accurate treatment outputs, marking them as strong baselines for further development.At the lower end, Gemini 2.0 Flash’s (77.4% ) score underscores how quickly performance can lag without iterative updates, while Deepseek-R1’s sub-80% result indicates potential weaknesses in domain-specific knowledge or recommendation specificity. Deepseek-chat’s solid upper-mid placement (81.9 %) shows that conversational optimization aids treatment advice, but it still falls short of models incorporating the newest architectural enhancements. Overall, these findings point to the importance of continuous model refinement to achieve clinically reliable treatment generation.
6.3 Treatment Follow-Up
Table 7: Follow up Results: Average Similarity Scores for Model-Generated Treatment Follow up Outputs.The underlined bold values are the best score.
Model Name Avg. Similarity score
Deepseek-R1 78.7
Deepseek-chat 79.4
Gemini 2.5 Pro Preview (03-25) 86.5
Gemini 2.0 Flash 66.5
Llama 3.3 70B 75.0
QWQ-32 83.7
Gemini 2.5 Flash Preview (04-17, Thinking) 82.6
Gemini 2.5 Flash Preview (04-17) 80.2
We generated model outputs using the Engineering template 3 and the Evaluation template 4 separately. As shown in Table 7, all eight models exhibited varied performance on generating treatment follow-up outputs, with average similarity scores ranging from (66.5 %) to (86.5 %). The Gemini 2.5 Pro Preview (03-25) led the group at (86.5 %), followed by QWQ-32 at (83.7 %) and the Gemini 2.5 Flash Preview (04-17, Thinking) at (82.6 %). The secondary Flash Preview (04-17) achieved (80.2 %), while Deepseek-chat and Deepseek-R1 produced (79.4 %) and (78.7 %), respectively. Llama 3.3 70B score (75.0 %), and the earliest variant, Gemini 2.0 Flash, trailed at (66.5 %). The substantial (20.0 %) spread between the top and bottom performers highlights that follow-up coherence poses a greater challenge than initial treatment or diagnosis tasks. Gemini 2.5 Pro Preview’s strong showing suggests its latest context-tracking optimizations significantly enhance continuity in multi-turn recommendations. QWQ-32’s robust (83.7%) score further underscores its ability to maintain patient context across successive exchanges. By contrast, Gemini 2.0 Flash’s steep drop to (66.5 %) indicates limitations in its reasoning architecture and narrower context window when managing follow-up prompts. The approximately (16.1 %) improvement from Gemini 2.0 Flash to the Gemini 2.5 Flash Preview (04-17, Thinking) illustrates the clear benefit of the "Thinking" inference mode for preserving treatment trajectories. Meanwhile, Deepseek-chat and Deepseek-R1, both below the (80.0% ) mark, demonstrate competence with basic follow-up queries but lack the advanced continuity afforded by newer model iterations. Llama 3.3 70B’s (75.0% ) result similarly suggests that even large foundational architectures may struggle with nuanced multi-turn follow-through without specialized enhancements.
6.4 Classification
Table 8: Classification Results: Disorder categories and specific disorders classification dataset evaluation on the chosen models. The underlined bold values are the best score for the specific column.
LLM Categories Disorder Specific Disorders
F1-Weighted Subset Accuracy F1-Weighted Subset Accuracy
DeepSeek R1 0.72 0.54 0.50 0.44
DeepSeek Chat 0.63 0.31 0.41 0.22
Gemini 2.5 Pro Preview (03-25) 0.60 0.26 0.42 0.23
Gemini 2.0 Flash 0.63 0.31 0.46 0.31
Llama 3.3 70B 0.68 0.46 0.47 0.34
QWQ 32 0.65 0.38 0.45 0.33
Gemini 2.5 Flash (04-17, Thinking) 0.60 0.26 0.37 0.14
Gemini 2.5 Flash (04-17) 0.68 0.47 0.44 0.36
We used the Engineering prompt template 5 to generate the model outputs, and then evaluated them for accuracy in classification without using a separate evaluation prompt.As shown in Table 8 compares each model’s ability to assign both broad disorder categories and granular specific disorders using two metrics: F1-weighted score and subset accuracy. Across the board, models achieve higher scores on the broader category task F1-weighted scores range from (0.60) to (0.72), while subset accuracy spans (0.26) to (0.54). In contrast, performance on specific disorders is uniformly lower, with F1-weighted values between (0.37) and (0.50) and subset accuracy from (0.14) to (0.44). This pattern reflects the greater difficulty of exactly matching the full set of fine-grained labels for each case.
DeepSeek-R1 leads in the category classification task, posting the highest F1-weighted score of (0.72) and subset accuracy of (0.54). Close behind, Llama 3.3 70B and the Gemini 2.5 Flash (04-17) variant both achieve balanced category performance, Llama at (0.68)/(0.46) and Gemini Flash at (0.68)/(0.47). For specific disorders, DeepSeek-R1 again tops the field with (0.50) F1-weighted and (0.44) subset accuracy, indicating its relative strength in identifying multiple, co-occurring disorders. QWQ-32 and Gemini 2.0 Flash also perform respectably on the specific task, with F1-weighted scores of (0.45)-(0.46).
The gap between category and specific-disorder scores, averaging around (0.20)-(0.25) F1 points, highlights the challenge of multi-label, fine-grained classification in clinical contexts. Models that excel at broad categorization still struggle to capture every nuanced diagnosis in a single pass, as evidenced by the lower subset accuracies across the board. These results underscore the need for targeted improvements, whether through additional domain data, hierarchical label modeling, or specialized prompt strategies to push LLMs toward clinically reliable, multi-label diagnostic performance.
6.5 Management Plan
Table 9: Management Plan Results: Average similarity scores of model-generated clinical management plans. The underlined bold values are the best score.
Model Name Avg. Similarity score
Deepseek-R1 84.4
Deepseek-chat 83.5
Gemini 2.5 Pro Preview (03-25) 84.1
Gemini 2.0 Flash 78.7
Llama 3.3 70B 82.1
QWQ-32 82.9
Gemini 2.5 Flash Preview (04-17, Thinking) 82.4
Gemini 2.5 Flash Preview (04-17) 81.1
We generate model outputs using the Engineering prompt template 6 and the Evaluation prompt template 7 separately.As shown in Table 9, all eight models produced high-quality management plans, with average similarity scores ranging from (78.7 %) to (84.4 %). This narrow spread of (5.7 %) points indicates that even the lowest-scoring model generates outputs that align reasonably well with expert-authored references, while the top performers exhibit only modest gains over the rest.DeepSeek-R1 achieved the highest score at (84.4%), marginally outperforming Gemini 2.5 Pro Preview (03-25) at (84.1 %) and DeepSeek-chat at (83.5 %). These leading models consistently synthesized comprehensive therapeutic recommendations, patient education points, and follow-up strategies that closely matched the reference plans. The strong performance of Gemini 2.5 Pro Preview underscores the impact of its recent architectural refinements on complex planning tasks. Mid-tier models QWQ-32 at (82.9 %), Gemini 2.5 Flash Preview (04-17, Thinking) at (82.4 %), and Llama 3.3 70B at (82.1 %) also deliver robust outputs, suggesting that both reasoning enhancements ("Thinking" mode) and broad contextual training contribute to solid plan generation. Gemini 2.5 Flash Preview (04-17) score (81.1 %), while Gemini 2.0 Flash lagged at (78.7 %), highlighting the clear benefit of iterative updates and instruction-tuning in newer model versions. Despite high similarity scores overall, careful human review remains essential to catch any subtle omissions or safety concerns before clinical use.
6.6 Clinical Approach
We generated model outputs using the Engineering prompt template 8 and the Evaluation prompt template 9 separately. As shown in Table 10 summarizes each model’s performance on the Clinical Approach task under two prompt formulations. Under the original prompt (Approach 1), average similarity scores range from (86.7 %) to (90.1 %), demonstrating that all models can reliably outline diagnostic frameworks and management plans when given a clear, familiar instruction. With the reformulated prompt (Approach 2), scores fall between (82.8 %) and (87.5 %), indicating a consistent sensitivity to prompt wording across architectures. Deepseek-R1 and Deepseek-chat start at the top of the leaderboard with (90.0 %) and (90.1 %) under Approach 1, respectively, but both experience moderate declines of (4.1 %) and (4.7 %) when switched to Approach 2. Llama 3.3 70B drops from (89.7 %) to (83.9 %) (- 5.8 %), while QWQ-32 falls from (89.2 %) to (82.8 %) (- 6.4 %), marking the largest sensitivity among the group. These results suggest that even high-capacity models can struggle to generalize reasoning steps if the prompt structure shifts.By contrast, Gemini 2.5 Pro Preview (03-25) exhibits the greatest robustness, declining only from (89.2 %) to (87.5 %) (- 1.7 %). Its Flash variants also remain relatively stable, with the "Thinking" mode decreasing by (2.1 %) (87.5 % → 85.4 %) and the standard Flash by (3.5 %) (88.6 % → 85.1 %). This relative consistency underscores the value of instructional tuning and specialized inference strategies in maintaining performance when prompt formulations vary.
Table 10: Clinical Approach Results: Average similarity scores for different language models across two prompt formulations on the Clinical Approach task. The "Change" column indicates the difference between Approach 2 and Approach 1. The underlined bold values are the best score for the specific column.
Model Name Avg. Similarity Score
Approach 1 Approach 2
Deepseek-R1 90.0 85.9
Deepseek-chat 90.1 85.4
Gemini 2.5 Pro Preview (03-25) 89.2 87.5
Gemini 2.0 Flash 86.7 85.1
Llama 3.3 70B 89.7 83.9
QWQ-32 89.2 82.8
Gemini 2.5 Flash Preview (04-17, Thinking) 87.5 85.4
Gemini 2.5 Flash Preview (04-17) 88.6 85.1
6.7 Mental QA
Table 11: Mental QA Results: Average similarity scores for different language models across two prompt versions on the Mental QA task. The "Change" column represents the difference between Prompt 1 and Prompt 2. The underlined bold values are the best score for the specific column.
Model Name Avg. Similarity Score
Prompt 1 Prompt 2
Deepseek-R1 88.0 88.1
Deepseek-chat 82.7 88.1
Gemini 2.5 Pro Preview (03-25) 87.0 84.6
Gemini 2.0 Flash 82.7 84.0
Llama 3.3 70B 84.2 84.1
QWQ-32 86.0 84.3
Gemini 2.5 Flash Preview (04-17, Thinking) 79.3 85.9
Gemini 2.5 Flash Preview (04-17) 83.5 83.1
We use the Engineering prompt template 10. As shown in Table 11, models’ average similarity scores under Prompt 1 range from (79.3 %) to (88.0 %), while under Prompt 2 they span (83.1 %) to (88.1 %). Deepseek-R1 remains near the top of both distributions with (88.0 %) and (88.1 %), respectively, and Deepseek-chat jumps from (82.7 %) to (88.1 %). At the lower end, the Gemini 2.5 Flash Preview (04-17, Thinking) rises from (79.3 %) to (85.9 %), whereas the Gemini 2.5 Pro Preview (03-25) declines from (87.0 %) to (84.6 %). These shifts underscore that definitional accuracy and its alignment with expert references varies substantially with prompt formulation.The magnitude of change varies by architecture. Deepseek-chat’s sizable +5.4 point gain and the Gemini 2.5 Flash (Thinking) model’s +6.6 point improvement highlight how enhanced instructional phrasing can bolster performance even in fact-based tasks. Conversely, the Gemini 2.5 Pro Preview’s -2.4 point drop demonstrates that a seemingly minor prompt tweak can grace alignment for some models. QWQ-32 (86.0 % → 84.3 %) and LLaMA 3.3 70B (84.2 % → 84.1 %) show minimal (<1 point) fluctuations, indicating greater prompt robustness in these cases.Deepseek-R1’s near-identical scores across both prompts suggest it generalizes definitional reasoning most consistently. Overall, these results emphasize the critical role of prompt Engineeringing in terminology-focused QA: small changes in phrasing or structure can yield double-digit performance swings. For sensitive domains like psychiatry, ensuring prompt clarity and close alignment with a model’s instruction-following tendencies is essential to achieve reliable, clinically accurate definitions.
6.8 Sequential Question Answering
Table 12: Sequential Question Answering Results: Average similarity scores for various language models on the Sequential Question Answering task using the baseline prompt (Prompt 1). The underlined bold values are the best score.
Model Name Avg. Similarity Score
Deepseek-R1 89.0
Deepseek-chat 88.9
Gemini 2.5 Pro Preview (03-25) 96.9
Gemini 2.0 Flash 90.0
Llama 3.3 70B 81.4
QWQ-32 86.2
Gemini 2.5 Flash Preview (04-17, Thinking) 89.7
Gemini 2.5 Flash Preview (04-17) 89.2
We generated model outputs using the Engineering prompt template 11 and the Evaluation prompt template12 separately.As shown in Table 12, all eight models exhibit strong performance on the Sequential Question Answering task under Prompt 1, with average similarity scores ranging from (81.4 %) to (96.9 %). Gemini 2.5 Pro Preview (03-25) stands out with an exceptional (96.9 %), demonstrating its superior ability to maintain context and deliver coherent, expert-aligned answers across multiple turns. Deepseek-R1 and Deepseek-chat follow closely at (89.0 %) and (88.9 %), respectively, indicating robust sequential reasoning capabilities.Mid-tier models including Gemini 2.5 Flash Preview (04-17, Thinking) at (89.7 %), Gemini 2.5 Flash Preview (04-17) at (89.2 %), and Gemini 2.0 Flash at (90.0 %) also deliver consistently high-quality outputs, showcasing effective context tracking and answer continuity. QWQ-32 achieves a solid (86.2 %), reinforcing its reliability in multi-turn clinical dialogues. These scores suggest that both "Thinking" inference modes and incremental flash updates materially enhance a model’s sequential QA performance. Llama 3.3 70B records the lowest similarity at (81.4 %), highlighting potential difficulties in capturing all nuanced dependencies between successive questions without additional prompt optimization. Nevertheless, its performance remains respectable, and prior analyses suggest that refined prompt Engineeringing could further close the gap. Overall, these results underscore that while top-tier proprietary models excel in sequential reasoning, prompt clarity and model updates play a critical role in elevating performance across the board.
6.9 Multiple Choice Questions (MCQ)
This section presents model performance on structured, objective assessment items commonly used in medical education. We start with single-best-answer questions (MCQs), followed by more complex matching and exam-style prompts. These formats test discrete factual recall, application of clinical reasoning, and the ability to integrate information across multiple elements.
6.9.1 Single-Answer Multiple-Choice Question
We used the Engineering prompt template 13 to generate the model outputs, and then evaluated them for accuracy in classification without using a separate evaluation prompt.As shown in Table 13, model performance on the MCQ task varies considerably depending on prompt style. Under the baseline prompt (P1), accuracy ranges from (75.5 %) to (83.2 %) and balanced accuracy from (64.3 %) to (79.3 %). With the few-shot prompt (P2), overall accuracy spans (71.6 %) to (82.8 %) and balanced accuracy (63.7 %) to (74.4 %). This indicates that while some models maintain consistency across both formats, others particularly those sensitive to prompt examples experience notable drops.Gemini 2.5 Pro Preview (03-25) delivers the most robust performance overall, scoring (81.7 %)/(79.3 %) in P1 accuracy/BA and improving to (82.8 %)/(74.4 %) in P2. Deepseek-R1 achieves the highest P1 accuracy at (83.2 %) but falls to (71.6 %) under P2, suggesting a strong grasp of standalone questions but shows better generalization when provided with examples. Llama 3.3 70B and QWQ-32 show stable results around (77.1 %)/(72.2 %) and (78.4 %)/(74.8 %) in P1, and (77.5 %)/(72.3 %) and (76.8 %)/(71.5 %) in P2 highlighting reliable reasoning unaffected by prompt format.By contrast, Gemini 2.0 Flash underperforms relative to its successors, with balanced accuracy at (64.3 %)/(65.2 %) in P1/P2 despite moderate accuracy (75.59 %)/(75.2 %). The Gemini 2.5 Flash Preview variants also show mixed results: the "Thinking" mode drops to (74.1 %)/(63.7 %) in P2, while the standard Flash holds (79.1 %)/(71.5 %). These findings underscore that MCQs though simpler than open-ended tasks still reveal differences in how models leverage prompt contexts and example-based cues to select correct answers.
Table 13: Multiple-choice question Results: Accuracy (A) and balanced accuracy (BA) scores for each model across all books and prompts.The underlined bold values are the best score for the specific column.
Model Name P1 P2
A BA A BA
Deepseek-R1 83.2 76.4 71.6 72.3
Deepseek-chat 75.8 71.5 74.4 69.1
Gemini 2.5 Pro Preview (03-25) 81.7 79.3 82.8 74.4
Gemini 2.0 Flash 75.5 64.3 75.2 65.2
Llama 3.3 70B 77.0 72.2 77.5 72.3
QWQ-32 78.4 74.8 76.8 71.5
Gemini 2.5 Flash Preview (04-17, Thinking) 79.4 69.8 74.1 63.75
Gemini 2.5 Flash Preview (04-17) 79.5 69.8 79.1 71.5
6.9.2 Extended Matching Item
We used the Engineering prompt template 13 to generate the model outputs, and then evaluated them for accuracy in classification without using a separate evaluation prompt. As shown in Table 14, all eight models achieve strong performance on the EMI task, with average similarity scores ranging from (72.2 %) to (87.94 %). Gemini 2.5 Pro Preview (03-25) leads the field under both prompts scoring (87.63 %) on Prompt 1 and (87.9 %) on Prompt 2 demonstrating exceptional consistency when matching complex, multi-option stems. Deepseek-R1 and Gemini 2.5 Flash Preview (04-17, Thinking) follow closely, both posting above (85.4 %) across disorderss, which underscores their robust reasoning capabilities even when instructions shift.In the mid tier, QWQ-32 records (78.5 %)/(79.8 %) under Prompts 1/2, while Deepseek-chat posts (80.34 %)/(79.8 %). These models handle extended contexts competently but display modest sensitivity to prompt wording, with score fluctuations under (1.1 %) points. Gemini 2.5 Flash Preview (04-17) shows the greatest prompt responsiveness, improving from (79.4 %) to (84.5 %) illustrating how optimized phrasing can yield substantial gains even on intricate matching tasks. At the lower end, both Gemini 2.0 Flash and Llama 3.3 70B trail their peers. Gemini 2.0 Flash drops from (80.1 %) on Prompt 1 to (75.5 %) on Prompt 2 a decline of (4.61 %) while Llama 3.3 70B remains flat at (72.3 %)/(72.2 %), indicating limited adaptation to the EMI format and instruction changes. These results highlight that, although modern LLMs can approach expert-level performance on EMIs, achieving maximum reliability requires careful prompt design and potentially EMI-specific instruction tuning. We also compared performance on the full-cluster versus separated-item EMI formats. As shown in Table 19, the full EMI format consistently outperforms the separated format, with average similarity scores of (84.1 %) versus (81.4 % )across all eight models. This finding further underscores the importance of preserving holistic context when matching complex, multi-option stems. The comparison between the full EMI format and the separated format is presented in section 10.2 in the appendix, where the full EMI format consistently outperforms the separated format across all eight models.
Table 14: Extended Matching Item Results: Average performance of different language models on the Extended Matching Item (EMI) task using full-question evaluation using two distinct prompts. The underlined bold values are the best score for the specific column.
Model Name Avg. Similarity Score
Prompt 1 Prompt 2
Deepseek-R1 85.5 85.8
Deepseek-chat 80.3 79.8
Gemini 2.5 Pro Preview (03-25) 87.6 87.9
Gemini 2.0 Flash 80.1 75.5
Llama 3.3 70B 72.3 72.2
QWQ-32 78.5 79.8
Gemini 2.5 Flash Preview (04-17, Thinking) 85.4 86.1
Gemini 2.5 Flash Preview (04-17) 79.4 84.5
6.9.3 Exams
Table 15: Exams Results: Average performance of different language models on the Exams task using full-question evaluation. The underlined bold values are the best score for the specific column.
Model Name Avg. Similarity Score
Prompt 1 Prompt 2
Deepseek-R1 79.2 74.3
Deepseek-chat 71.4 69.2
Gemini 2.5 Pro Preview (03-25) 80.4 71.3
Gemini 2.0 Flash 69.9 71.1
Llama 3.3 70B 66.5 65.7
QWQ-32 73.4 67.2
Gemini 2.5 Flash Preview (04-17, Thinking) 78.4 76.8
Gemini 2.5 Flash Preview (04-17) 73.1 71.6
We used the Engineering prompt template 13 to generate the model outputs, and then evaluated them for accuracy in classification without using a separate evaluation prompt.As shown in Table 15, model performance on the Exams task under Prompt 1 ranges from (66.5 %) to (80.4 %), while under Prompt 2 it spans (65.7 %) to (76.8 %). Gemini 2.5 Pro Preview (03-25) leads the cohort on Prompt 1 with (80.4 %), reflecting its strong ability to comprehend and synthesize multi-sentence clinical vignettes into expert-aligned answers. However, its drop to (71.3 %) under Prompt 2 highlights that its exam-style reasoning is sensitive to instruction phrasing. Deepseek-R1 posts the next highest scores at (79.2 %)/(74.3 %), indicating reliable baseline performance and moderate resilience to prompt reformulation. Gemini 2.5 Flash Preview (04-17, Thinking) variant demonstrates the greatest robustness, scoring (78.4 %) on Prompt 1 and maintaining a high (76.8 %) on Prompt 2 a decline of only (1.68 %) suggesting that its enhanced reasoning mechanisms effectively preserve context and structure even when instructions change. Mid-tier models such as QWQ-32 ((73.4 %)/(67.2 %)), Gemini 2.5 Flash Preview (04-17) at ((73.1 %)/(71.6 %)), and Deepseek-chat ((71.4 %)/(69.2 %)) show competent understanding of exam questions but experience steeper drops of (5-6 %) under the alternate prompt, implying they rely more heavily on specific cue wording.At the lower end, Gemini 2.0 Flash and Llama 3.3 70B struggle the most: Gemini 2.0 Flash edges up slightly from (69.9 %) to (71.1 %) when given the second prompt perhaps benefiting from clearer examples while Llama 3.3 70B remains nearly flat at (66.5 %)/(65.7 %). These results suggest that without targeted prompt Engineeringing or exam-specific tuning, foundational models may underperform on complex, structured assessment items. Overall, the data underscore that both advanced model architectures and prompt clarity are critical for reliable automated evaluation of professional-level exam questions.
6.10 Benchmarking LLMs Against Top Performers
Table 16: Performance Comparison of Various LLMs on Clinical and Mental-Health Tasks
Task MedGemma Sonnet 4 QWEN 3 32 (Score) Best Model
Diagnoses 79.4 85.7 82.1 (86%) Gemini 2.5 Pro
Treatment 84.2 84.5 81.6 (84.5%) Sonnet 4
Treatment Follow-Up 79.2 83.5 80.9 (86.5%) Gemini 2.5 Pro
Classification (Disorder Categories) 0.12, 0.46 0.26 (0.54%) DeepSeek R1
Classification (Specific Disorders) 0.03 0.34 0.14 (0.44%) DeepSeek R1
Management Plan 81.7 84.8 80.3 (84.8%) Sonnet 4
Clinical Approach 90.0 90.0 90.0 (90.01%)Deepseek chat
Mental QA 87.1 89.35 90.6 (90.6%) QWEN 3 32
Sequential QA 87.7 89.6 89.7 (96.5%) Gemini 2.5 Pro
MCQ 87.4 81.6 74.2 (87.4%) MedGemma
Exam 69.2 77.0 65.3 (80.4%) Gemini 2.5 Pro
EMI 72.5 83.9 73.1 (87.9%) Gemini 2.5 Pro
Table 16 summarizes the performance of four models, including MedGemma, Sonnet 4, QWEN 3 32B, and Gemini 2.5 Pro across eleven clinical and mental-health tasks. Overall, Sonnet 4 and Gemini 2.5 Pro achieve the highest scores on the majority of tasks. Sonnet 4 leads on Treatment (84.5 %), Management Plan (84.8 %), and Mental QA (89.35 %), while Gemini 2.5 Pro outperforms on Diagnoses (86 %), Treatment Follow-Up (86.5 %), Sequential QA (96.5 %), Exam (80.4 %), and EMI (87.9 %). Deepseek-chat narrowly tops Clinical Approach at 90.01 %, essentially tying with the others at 90 %. DeepSeek R1 demonstrates strength in broad Classification (Disorder Categories) (A = 0.54, F1 = 0.72), and on Specific Disorders Classification (A = 0.44, F1 = 0.50).No single model dominates across all tasks. Gemini 2.5 Pro excels in sequential reasoning and recall-intensive tasks such as Sequential QA and Exam, likely due to its larger parameter count and advanced contextual understanding. QWEN 3 32B strong performance in treatment-related tasks and mental QA suggests effective fine-tuning on healthcare dialogues. MedGemma’s superior classification of specific disorders indicates strength in entity recognition and fine-grained labeling, while QWEN 3 32B’s relative success on broader disorder categorization may reflect robust generalization.The modest F1 scores in classification tasks highlight an area for improvement: models struggle with balanced precision and recall when differentiating nuanced categories. Future work should explore ensemble approaches or targeted fine-tuning on underrepresented classes to boost classification F1. Additionally, incorporating external medical knowledge bases may further enhance diagnostic and treatment reasoning.
6.11 Comparative Performance Radar of All Models
This radar chart presents the comparative performance of 12 LLMs across eleven psychiatric reasoning tasks from the PsychiatryBench. Each axis corresponds to a distinct task category, including Diagnosis, Treatment, Treatment Follow-Up, Classification (Categories and Specific Disorders), Management Plan, Clinical Approach, Mental QA, Sequential QA, MCQ, and EMI. The scores reflect similarity-based evaluation metrics computed using an LLM-as-a-judge method with Llama 3 70B as the evaluator. Notably, Gemini 2.5 Pro, Gemini 2.5 Flash Thinking, and QWQ-32 consistently perform near the top across most domains, particularly in Diagnosis, Sequential QA, and EMI, while Deepseek R1 and Qwen 3 32 show relatively lower performance in Classification tasks. The pronounced dip in scores for "Classification (Specific Disorders)" highlights a consistent challenge across all models. This visualization underscores the strengths and limitations of each model across complex psychiatric tasks, providing insights into areas requiring further model refinement.
Figure 3: Radar chart showing LLM performance across $1$ psychiatric tasks in the PsychiatryBench. Models like Gemini 2.5 Pro and QWQ-32 lead overall, while classification tasks remain the most challenging.
7 Limitations and Future Work
Although this study offers valuable information on the capabilities of LLMs for mental health tasks, it is important to recognize several limitations that may affect the interpretation and generalizability of our findings. In addition to the future directions that can help achieve higher performance.
7.1 Limitations
While PsychiatryBench represents a significant step towards the clinically grounded evaluation of LLMs in psychiatry, its development and scope were shaped by several inherent limitations. These constraints, which our work seeks to begin addressing, stem primarily from the nature of the available source material and the practicalities of evaluation at scale.
The foremost limitation is a direct consequence of the critical research gap we identify: the absence of large-scale, publicly available, and expert-validated clinical datasets for mental health. Forced a reliance on meticulous manual curation from authoritative textbooks. While this process ensures unparalleled clinical validity, the core contribution of our work is inherently labor-intensive. As a result, the scale of PsychiatryBench, though substantial, is smaller than benchmarks built on synthetic data or automated web-scraping. It was a deliberate trade-off to prioritize clinical safety over sheer data volume.
Furthermore, our reliance on published textbooks introduces a trade-off between authoritative knowledge and real-world clinical practice. The vignettes in these resources are, by design, well-structured and pedagogically focused to illustrate specific diagnostic principles. They may not fully capture the ambiguity, noise, and unstructured nature of raw clinical data, such as real-time patient dialogues or fragmented electronic health record (EHR) notes. Consequently, while PsychiatryBench robustly tests an LLM’s reasoning against expert-curated knowledge, performance may not perfectly generalize to the complexities of live, unfiltered clinical interactions.
A related point is that the distribution of psychiatric disorders within PsychiatryBench does not mirror their real-world epidemiological prevalence. It is a direct consequence of our source materials; psychiatric textbooks often over-represent rare, complex, or "classic" cases to serve specific pedagogical goals. Therefore, models evaluated on PsychiatryBench are tested on their ability to handle a teaching-oriented distribution of cases, which differs from the case mix they would encounter in a general clinical setting.
Finally, our evaluation methodology for open-ended tasks relies on LLM-as-a-judge. While this approach offers the scalability and consistency required to evaluate thousands of responses across numerous models, it is not a perfect substitute for human expert evaluation. LLM judges can have inherent biases and may not capture the full clinical nuance apparent to a practicing psychiatrist. We mitigated this by grounding the evaluation in expert-validated reference answers, meaning the judge’s task was one of comparing an output to a "gold standard" rather than making a de novo clinical judgment. However, a full-scale validation by a panel of human psychiatrists was beyond the scope of this work and remains a crucial future direction.
These limitations highlight the very challenges our work was designed to confront. The constraints on scale, data distribution, and evaluation methods are direct consequences of the field’s lack of high-fidelity psychiatric benchmarks. By navigating these issues and creating a transparent, clinically-grounded resource, PsychiatryBench establishes an essential baseline. It serves not only as an evaluation tool but as a foundational scaffold upon which future research can build to address these very limitations, paving the way for safer and more reliable AI in mental healthcare.
7.2 Future Work
Our analysis identifies seven key priorities for advancing LLM applications in mental health, summarized in the following points:
  • Expanded Dataset Coverage: Future research should aim to include a wider range of psychiatric disorders, cultural contexts, and multimodal data sources (e.g., EHRs, imaging, and audio transcripts) to create a more representative benchmark. Collaborations with international psychiatric organizations could facilitate access to diverse clinical data.
  • Automated and Scalable Annotation: Developing robust automated tools for case extraction and annotation, validated by expert psychiatrists, would enable larger-scale datasets while maintaining clinical accuracy. Techniques like natural language processing (NLP) for structured data extraction could streamline this process.
  • Multimodal Evaluation Frameworks: Integrating multimodal inputs, such as lab results or patient audio, into evaluation datasets would better simulate real-world clinical decision-making. Could involve adapting LLMs to process structured data alongside text or developing hybrid models for multimodal reasoning.
  • Broader LLM Comparisons: Evaluating a diverse set of LLMs, including open-source models and those fine-tuned for medical domains, would provide deeper insights into their relative strengths and weaknesses in psychiatric applications.
  • Enhanced Evaluation Metrics: Future work should focus on designing metrics that prioritize clinical safety, cultural sensitivity, and robustness against hallucinations. Incorporating human-in-the-loop evaluation by practicing psychiatrists could ensure alignment with clinical standards.
  • Comprehensive Task Diversity: Expanding the task set to include empathetic dialogue, crisis intervention, and longitudinal care planning would enable a more complete assessment of LLMs’ potential in mental healthcare. Could involve creating conversational benchmarks like MentalChat 16K but grounded in expert-validated data.
  • Prompt Optimization Frameworks: Systematic exploration of the prompt space, using automated prompt engineering tools, could identify optimal prompt structures for clinical tasks. Standardizing prompt formats across studies would improve reproducibility and comparability of results.
8 Conclusion
In this work, we have introduced PsychiatryBench, the first comprehensive, clinically grounded benchmark for evaluating LLMs in psychiatric applications. By curating over 5,300 question-answer pairs from eleven authoritative textbooks and casebooks, ranging from diagnostic vignettes and treatment planning to longitudinal follow-up and complex sequential reasoning, we ensure that model performance is measured against expert-validated, real-world clinical knowledge rather than unverified or synthetic data. Our eleven distinct task types probe the full spectrum of psychiatric reasoning, from factual recall and classification to multi-turn decision-making and management planning.
Through rigorous evaluation of both general-purpose and medical-specialized LLMs, including proprietary systems such as Gemini 2.5 Pro and MedGemma, as well as open-source models like LLaMA and Palmyra-Med, we identify clear performance leaders and persistent gaps. Frontier models achieve strong alignment on single-turn tasks such as diagnosis and management planning (up to 8̃7 % similarity). However, performance degrades on multi-step follow-up, classification of comorbid disorders, and prompt robustness across varying formats. Our novel similarity scoring, along with consistency metrics for full versus separated EMIs and multi-label classification, provides a nuanced picture of models’ strengths and limitations in high-stakes clinical assessments.
These findings underscore both the promise and the perils of deploying LLMs in mental healthcare. On one hand, leading models demonstrate impressive capacity for clinical reasoning, empathetic response, and knowledge synthesis; on the other, the risk of hallucinations, prompt sensitivity, and difficulties with multi-label, longitudinal scenarios highlight the necessity of human oversight and continued refinement.
Looking ahead, PsychiatryBench offers a modular framework for future research. Extending the dataset with additional case complexities, such as cross-cultural presentations, pediatric and forensic psychiatry, and integrating multimodal inputs (e.g., speech, imaging) will further bridge the gap between AI evaluation and real-world clinical practice. Moreover, fine-tuning and retrieval-augmented approaches hold promise for improving model robustness and safety. We encourage the community to leverage PsychiatryBench as both a diagnostic tool to benchmark new models and as a scaffolding for developing clinically reliable, interpretable AI systems that can augment psychiatric care while rigorously safeguarding patient well-being.
9 Ethical Statement, Copyright, and Fair Use Statement
The PsychiatryBench dataset was developed using content manually curated from publicly available, expert-authored psychiatric textbooks and casebooks. All material was selected and annotated by subject matter experts with a clinical background to support research in safe and effective psychiatric reasoning by AI models. No personal, identifiable, or patient-derived data were used in the construction of this dataset. The dataset does not contain any clinical notes, transcripts, or sensitive medical information from real individuals.
This research adheres to ethical guidelines regarding the use of educational content for research purposes. The dataset is intended solely for non-commercial research, academic, and educational uses that aim to improve the safety, reliability, and clinical alignment of language models in mental health applications. All references to psychiatric diagnoses, treatments, and case scenarios are drawn from simulated or publicly available educational materials, not real patient interactions.
All content in the PsychiatryBench dataset is derived from publicly available and widely used psychiatric educational resources for academic research and benchmarking. The inclusion of questions and scenarios from books such as DSM-5-TR Clinical Cases, Stahl’s Psychopharmacology, and others is done under the principle of fair use, specifically for:
  • Transformative use: The material has been substantially modified, structured, and reformatted into a new dataset format for the distinct purpose of evaluating AI systems.
  • Non-commercial research: The dataset is intended solely for research and academic purposes, without any commercial exploitation.
  • Limited scope and effect on market: Only selected, representative samples were used, not the entirety of any single source, and use of the dataset does not substitute for or compete with the original textbooks.
Users of the dataset are responsible for ensuring compliance with local copyright laws. They should obtain appropriate permissions if intending to reuse substantial parts of the original material for commercial or distribution purposes.
Supporting Information: Does the "Candidate Answer" include the key symptoms, test results,
or other evidence that support the diagnoses? It doesn’t need to include *every detail, but
it should capture the most important ones. It’s ok if this supporting information is
implicitly present in the reference answer.
Exclusion of Irrelevant Information: Do *not penalize the "Candidate Answer" for including
additional information that is not present in the "Reference Answer", *as long as that
information is medically accurate and relevant to the case*.
groundance does not affect the score: Do not penalize a response for being redundant.
4. Provide a single numerical similarity score between 0 and 100, where:
0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms of
    diagnosis or supporting information. It’s completely wrong.
50: The "Candidate Answer" identifies *some relevant diagnoses or supporting information,
    but it misses major elements or includes significant inaccuracies.
75: The "Candidate Answer" correctly identifies the primary diagnosis (if applicable) and
    most of the important differential diagnoses and supporting information. It may miss some
    minor details or include some slight inaccuracies.
90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core diagnostic
    meaning and completeness. It may have minor omissions or differences in emphasis, but it
    captures almost all of the essential information.
100: The "Candidate Answer" is a perfect match in terms of core diagnostic meaning. It
    includes all of the key diagnoses and supporting information from the "Reference Answer",
    even if the wording or organization is different. Do not justify or explain output only
    the numeric score.
5. Do not justify or explain the answer. Give a score between 0 and 100, no other
text. Now, evaluate the following
Concept to Define : {Answer}
Candidate Response: {Response}
Template 3: Treatment Follow Up Prediction Prompt
As a psychiatrist, analyze the provided patient’s information, history, and follow ups to
answer the question according to the guidelines.
Guidelines
- Provide reasoning in your answer.
- Don’t use bullet points just sentences.
- No introduction before the answer just provide the answer for the provided question only.
The Patient’s History: {History}
if {First Follow Up}:
    follow = f"""First Follow Up
    {First Follow Up}
    prompt_template = prompt_template + follow
if {Second Follow up}:
    follow = f"""Second Follow up:
    {Second Follow up}
    prompt_template = prompt_template + follow
q = f"""Question: {Question}"""
prompt_template = prompt_template + q
Template 4: Follow Up Evaluation Prompt
SYSTEM: You are a mental health case response evaluator. Your task is to compare two text
responses about a specific mental health case, assessing the degree to which the first text
(the "Candidate Answer") captures the essential information regarding the case assessment and
recommended follow-up contained in the second text (the "Reference Answer"). Focus on the core
clinical meaning, the completeness of the assessment understanding, and the
appropriateness/completeness of the follow-up plan, *not on writing style or specific wording.
User:
Instructions:
    1. Read the "Reference Answer" carefully. Identify the key aspects of the case assessment
    (e.g., presenting problem, relevant history, potential diagnoses/formulations, risk
    assessment) and the core elements of the recommended follow-up plan (e.g., therapy
    recommendations, medication considerations, further assessments needed, safety planning,
    referrals).
    2. Read the "Candidate Answer". This answer will also describe an assessment of the case
    and propose a follow-up plan.
    3. Compare the "Candidate Answer" to the "Reference Answer" based on the following criteria:
        Assessment Accuracy & Completeness: Does the "Candidate Answer" accurately reflect the
        core understanding of the patient’s situation (presenting issues, key historical factors,
        potential clinical formulation/diagnosis) as presented in the "Reference Answer"? Does
        it capture the most critical elements of the assessment?
        Follow-up Plan Completeness & Appropriateness: Does the "Candidate Answer" include the
        essential elements and intent of the follow-up plan outlined in the "Reference Answer"?
        This includes key interventions (therapy, medication), monitoring, safety considerations,
        and necessary referrals.Minor variations in terminology are acceptable if the underlying
        clinical concept and action are the same.
        Rationale/Supporting Information: Does the "Candidate Answer" implicitly or explicitly
        capture the key reasons or clinical justifications for the assessment and follow-up plan
        that are present in the "Reference Answer"? It doesn’t need every detail but should align
        on the core reasoning.
        Exclusion of Irrelevant Information: Do not penalize the "Candidate Answer" for including
        additional information *as long as that information is clinically accurate and relevant
        to the mental health case and its management*. However, significantly contradictory or
        inappropriate additions may lower the score.
        groundance does not affect the score: Do not penalize a response for being
        redundant.
    4. Provide a single numerical similarity score between 0 and 100, where:
        0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms
        of case assessment or follow-up plan. It’s completely different or clinically
        inappropriate.
        50: The "Candidate Answer" identifies some relevant aspects of the assessment or follow-
        up plan but misses major elements, contains significant inaccuracies, or proposes a
        substantially different approach.
        75: The "Candidate Answer" correctly captures the main aspects of the assessment and most
        of the essential follow-up recommendations from the "Reference Answer". It may miss some
        minor details, nuances, or justifications.
        90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core case
        understanding and the proposed follow-up plan. It may have minor omissions or differences
        in emphasis but captures almost all essential clinical information and actions.
        100: The "Candidate Answer" perfectly matches the "Reference Answer" in terms of core
        clinical meaning regarding the assessment and follow-up plan. It includes allkey
        assessment points and recommended actions, even if the wording or organization differs.
        Do not justify or explain the answer. Give a score between 0 and 100, no other text.
Now, evaluate the following
Reference Answer: {Answer}
Candidate Response: {Response}
Template 5: Classification Prompt
    As a psychiatrist, analyze the provided patient’s information and history to determine if
    the patient exhibits clear symptoms of any of the provided list of disorders according to
    provided guidelines.
    Guidelines
    - The patient could have multiple illness at the same time else he is normal.
    - Concise Response: Respond with any combination of these disorders separated by a comma
    and if the poster doesn’t have any mental illness just answer with Normal.
    - No Explanations: Don’t providing explanations for your assessment.
    - Ambiguity: If the it is unclear, choose the most probable label/disorder.
    Disorders List:
    The Disorders List (Categories or Specific Disorders)
    The Patient’s History:
    {History}
Template 6: Management Plan Prediction Prompt
You are a psychiatrist. Analyze the provided patient’s information(History, Mental State
Examination, and Physical Examination) to answer the psychiatric protocol question according
to the guidelines below.
Guidelines:
- Provide clear reasoning in your answer based on the information provided.
- Use full, coherent sentences only; do not use bullet points.
- Do not include an introduction or any prefatory remarks; start directly with the answer.
Patient’s History:
{History}
Patient’s Mental State Examination:
{Mental State Examination}
Patient’s Physical Examination:
{Physical Examination}
Question:
{Question}
Template 7: Management Plan Evaluation Prompt
SYSTEM: You are a psychiatric protocol response evaluator. Your task is to compare two text
responses about a psychiatric protocol question, assessing the degree to which the first text
(the "Candidate Answer") captures the essential information contained in the second text (the
"Reference Answer" or "True Label"). Focus on the clinical correctness, completeness of
[…]
User:
Instructions:
1. Read the "Reference Answer" carefully. Identify the key components expected in the
    psychiatric protocol response (e.g., symptom recognition, diagnostic steps, risk/safety
    considerations, treatment recommendations, ethical/legal procedures if applicable).
2. Read the "Candidate Answer". This answer will attempt to respond to the same psychiatric
    protocol question.
3. Compare the "Candidate Answer" to the "Reference Answer" based on the following criteria:
     Protocol Adherence & Accuracy: Does the "Candidate Answer" correctly follow the
    psychiatric protocol steps presented in the "Reference Answer"? Does it recognize key
    clinical issues, decision points, and require actions?
     Completeness of Response: Does the "Candidate Answer" address all major elements expected
    by the protocol as reflected in the "Reference Answer" (e.g., diagnostic considerations,
    intervention steps, risk management)?
     Reasoning and Justification Alignment: Does the "Candidate Answer" capture the rationale
    (explicit or implied) behind the steps taken according to the protocol in the "Reference
    Answer"? It should reflect the clinical reasoning that justifies assessment and management
    decisions.
     Handling of Additional Information: Extra information is acceptable if it is clinically
    correct and protocol-consistent. Do not penalize additional relevant detail. However,
    incorrect, contradictory, or unsafe suggestions should lower the score. groundance
    Tolerance: groundance (repeating points) does not affect the score.
4. Provide a single numerical similarity score between 0 and 100, where:
     0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms
    of protocol steps, clinical reasoning, or appropriate action. 50: The "Candidate Answer"
    captures some relevant elements but misses major steps, has serious inaccuracies, or
    suggests significantly incorrect approaches. 75: The "Candidate Answer" captures most key
    protocol elements and clinical reasoning, with only minor omissions or inaccuracies.
     90: The "Candidate Answer" is very close to the "Reference Answer" in terms of
    psychiatric protocol adherence, clinical assessment, and recommended actions, with only
    slight gaps or minor emphasis differences. 100: The "Candidate Answer" fully matches the
    "Reference Answer" in terms of psychiatric protocol steps, reasoning, and recommended
    actions, even if wording or order differ.
5. Do not justify or explain the answer. Only provide a single numerical score between 0 and
100. No other text output.Now, evaluate the following
Reference Answer: {Answer}
Candidate Response: {Response}
Template 8: Clinical approach Prompt
Patient History: {History}
Clinical Approach: {Clinical Approach}
Generated Question: Generate exactly one clear, focused, and medically relevant question
based on Clinical Approach.
Clinical approach Question Generation Prompt
You are an expert grader evaluating whether a history question is relevant to a given
clinical approach.
Question: {Question_Generation}
Clinical Approach: {Clinical Approach}
Determine if the question contains keywords or semantic meaning that are directly related to
the clinical approach. Respond only with "yes" if it is relevant or "no" if it is not. Do
not provide any explanation.
Clinical approach Grading Prompt
You are a psychiatrist. Based on the patient’s clinical history and the
question, provide a clear, accurate, and medically appropriate answer.
Patient History:{History}
Question: {Question_Generation}
Clinical approach Answer Generation Prompt
You are a clinical reasoning expert. Based on the following patient case history, write a
concise summary of the clinical approach. Include possible diagnoses, recommended
investigations, and management steps. The tone should be similar to a medical case
explanation, like those found in exam review books.
Patient History:{History}
Write a clinical approach (1-2 paragraphs):
Clinical Approach Prediction Prompt
Template 9: Clinical Approach Evaluation Prompt
SYSTEM: You are a mental health case response evaluator. Your task is to compare two text
responses about a specific mental health case, assessing the degree to which the first text
(the "Candidate Answer") captures the essential information regarding the clinical approach
contained in the second text (the "Reference Answer"). Focus on the clinical reasoning,
assessment strategy, and diagnostic insight not on writing style or specific wording.
User:
Instructions:
1. Read the "Reference Answer" carefully. Identify the key components of the clinical
    approach: presenting symptoms, history taking, differential diagnosis, initial
    evaluation steps, and rationale for assessment strategy.
2. Read the "Candidate Answer". This is another attempt to outline a clinical approach to
    the same case.
3. Compare the "Candidate Answer" to the "Reference Answer" based on the following criteria:
    Assessment Strategy Accuracy & Completeness: Does the "Candidate Answer" reflect an
    accurate and complete approach toassessing the case?
    Does it capture the key symptoms, relevant history, and potential clinical hypotheses
    found in the "Reference Answer"?
    Diagnostic Reasoning & Thought Process: Does the "Candidate Answer" follow a similar
    line of clinical reasoning, including formulation or differential diagnosis as in the
    "Reference Answer"?
    Clinical Actions or Evaluation Steps: Does it propose similar or equivalent steps in the
    evaluation process (e.g., screeningtools, physical exams, lab tests, psychiatric
    interview elements)?
    Rationale Alignment: Does the "Candidate Answer" align with the clinical logic and
    justifications provided in the "Reference Answer"? Perfect wording is not needed focus
    on reasoning.
    No Penalty for Relevance: Do not penalize clinically valid additions if they are
    appropriate, even if not found in the"Reference Answer". groundance does not grace score.
4. Provide a single numeric score between 0 and 100:
0: Completely off-topic, unrelated, or clinically incorrect.
50:Some valid elements of clinical approach are present, but many key elements are missing or
poorly reasoned.
75: Mostly aligned with correct steps and rationale, missing only some secondary elements.
90: Nearly identical clinical reasoning and plan; may differ in phrasing.
100: Fully matches the core clinical approach and rationale.
Do not justify or explain output only the numeric score.
Now, evaluate the following
Reference Answer: {clinical approach}
Candidate Response: {Response}
Template 10: Three-stage Mental QA prompt pipeline. (a) Prompt 1 directs the model to answer clinical questions step-by-step using patient history in a structured, professional tone; (b) Prompt 2 refocuses on defining the target mental-health concept clearly and concisely, as in exam-style review; (c) Evaluation Prompt instructs an automated scorer to assign a 0-100 accuracy score based on completeness, clarity, depth, and relevance of the model’s definition.
You are a psychiatrist specializing in mental health knowledge. Based on the provided dataset,
answer the following question accurately
and concisely. Ensure your response is well-structured, using bullet points or numbered lists
where appropriate, and maintain a professional
tone throughout. Think step by step to provide a logical and coherent answer like those found
in exam review books.
History: {History}
Question: {Question}
Mental QA Prediction Prompt 1
You are an expert psychiatrist with deep knowledge of mental health concepts. Your task is to
use the patient’s history to answer the
clinical question by defining or explaining the target concept clearly and concisely like
those found in exam review books.
History: {History}
Question: {Question}
Mental QA Prediction Prompt 2
System: You are a mental health concept knowledge evaluator. Your task is to assess how
accurately, completely, and clearly the candi-
date’s response defines the concept provided in the "Answer" field, taking into account the
clinical context in the "History."
User:
Instructions:
    1. Read the "Answer" this is the clinical concept or term to define (e.g., "Loss of
    interest or pleasure in activities. . . ").
    2. Read the "Candidate Response" the model’s definition/explanation of that concept.
    3. Evaluate the response on:
    Definition Accuracy & Completeness: Are all core features of the concept present and
    correctly described?
    Clarity & Precision: Is the explanation clear, unambiguous, and clinically precise?
    Depth of Explanation: Does it include relevant examples or elaborations that demonstrate
    understanding?
    Relevance & Focus:Does it avoid irrelevant details and stick to the concept at hand?
    4. Provide a single numeric score between 0 and 100:
    0: No meaningful overlap incorrect or missing core elements.
    50: Some correct elements but major omissions or inaccuracies.
    75: Mostly correct with only minor gaps or imprecisions.
    90:Very close to a perfect definition; only small details missing.
    100:Perfectly accurate, complete, and clear.
    Do not justify or explain output only the numeric score.
Now, evaluate the following
Concept to Define : {Answer}
Candidate Response: {Response}
Mental QA Evaluation Prompt
Template 11: Sequential Question Answering Prediction Prompt
Patient History: {History}
Task: Based on the provided history, answer the following five clinical questions. Use the
same format and level of detail shown in the example. Make sure to include preferred and
alternative diagnoses where applicable, and support your answers with reasoning drawn fromthe
case.
Q1: What is the likely differential diagnosis?
    preferred diagnosis:
    Alternative diagnoses:
Q2: What information in the history supports the diagnosis, and what other information would
    help to confirm it?
    (Provide a detailed rationale using history details and explain what additional
    information/tests are needed to confirm the diagnosis.)
Q3 : What might the important etiological factors be?
    (Organize into: Pregrazing, Precipitating, Maintaining factors)
Q4: What treatment options are available?
    (Cover location-based options, physical treatments, psychological support, and social
    considerations. Mention use of Mental Health Act if relevant.)
Q5: What is the prognosis in this case?
    (Discuss short-term and long-term outlook. Include good and poor prognostic indicators
    based on patient specifics.)
Q6: based on the provided History ,which class does this diagnosis belong to according to
    DSM-5 ?
Template 12: Sequential Question Answering Evaluation Prompt
SYSTEM: You are a mental health case response evaluator. Your task is to assess six clinical
responses from a trainee clinician. Each response corresponds to a specific psychiatric
question.
For each question, compare the candidate’s answer to the expert reference answer and provide a
numeric score based on how well the candidate captures the clinical reasoning and content.
    • Assign a similarity score between 0 and 100:
        - 0: Text A does not contain the correct diagnoses or significantly misrepresents key
        clinical reasoning compared to Text B.
        - 50: Text A identifies some key diagnoses but either misses important differentials or
        partially misrepresents clinical reasoning compared to Text B.
        - 100 means Text A correctly captures all essential diagnoses, differential diagnoses,
        and clinical reasoning as stated in Text B.
    • Use the following criteria to guide your scoring:
        - Presence of the correct main diagnosis (40 points).
        - Inclusion and accuracy of differential diagnoses (30 points).
        - Accuracy and completeness of clinical reasoning and justification (30 points).
    • Provide ONLY the numeric similarity score (0-100) without any additional explanation or
    justification.
Template 13: Multi-stage prompting strategies for objective assessment items. (a) MCQ Prompt 1 directs the model to choose one answer without explanation; (b) MCQ Prompt 2 adds a brief analytical step over patient history; (c) EMI Prompt 1 handles extended matching items where options may be reused or omitted; (d) EMI Prompt 2 blendsn-based EMIs with standard MCQs, requiring only the answer letters.
Answer the question using one of the given choices.
Do not provide any explanation for your choice. Indicate your choice with a single English
letter. (Provide only one answer; Multiple answers could be provided.)
History:{History}
Question:{Question}
Choices:{Choices}
MCQ Prompt 1
Analyze the multiple-choice question (MCQ) and select the best answer.
Assess each option logically based on relevant psychological history.
Do not provide any explanation for your choice. Indicate your choice with a single English
letter. (Provide only one answer; Multiple answers could be provided.)
History:{History}
Question:{Question}
Choices:{Choices}
MCQ Prompt 2
Analyze the extended matching items (EMI) and select the best answer. each option (denoted a,
b, c, etc.)
might be used once, more than once, or not at all. Assess each option logically. Please don’t
give any explanation for your choice.
Question:{Question}
Choices:{Choices}
EMI Prompt 1
Analyze the multiple-choice question (MCQ) and select the best answer.
Assess each option logically.
Do not provide any explanation for your choice. Indicate your choice with a single English
letter. (Provide only one answer; Multiple answers could be provided.)
EMI Prompt 2
Template 14: Point-based similarity EMI prompting templates: (a) a straightforward "header-required" template that simply presents the fields, (b) an analytical variant that adds an instruction to assess each option logically before choosing.
Answer the question using one of the given choices.
Do not provide any explanation for your choice. Indicate your choice with a single English
letter. (Provide only one answer; Multiple answers could be provided.)
Header:{Header}
Required:{Required}
Question:{Question}
Choices:{Choices}
(a) EMI Separated Prompt 1
Analyze the multiple-choice question (MCQ) and select the best answer.
Assess each option logically .
Do not provide any explanation for your choice. Indicate your choice with a single English
letter. (Provide only one answer; Multiple answers could be provided.)
Header:{Header}
Required:{Required}
Question:{Question}
Choices:{Choices}
(b) EMI Separated Prompt 2
Template 15: Prompt generated by GPT-4.5 for point-based similarity scoring of clinical reasoning responses.
I will provide two texts: Text A (LLM’s response) and Text B (True answer). Evaluate their
similarity based strictly on the presence and accuracy of key diagnoses, differential
diagnoses, and clinical reasoning. Follow these guidelines:
    • Assign a similarity score between 0 and 100:
        - 0: Text A does not contain the correct diagnoses or significantly misrepresents key
        clinical reasoning compared to Text B.
        - 50: Text A identifies some key diagnoses but either misses important differentials or
        partially misrepresents clinical reasoning compared to Text B.
        - 100 means Text A correctly captures all essential diagnoses, differential diagnoses,
        and clinical reasoning as stated in Text B.
    • Use the following criteria to guide your scoring:
        - Presence of the correct main diagnosis (40 points).
        - Inclusion and accuracy of differential diagnoses (30 points).
        - Accuracy and completeness of clinical reasoning and justification (30 points).
    • Provide ONLY the numeric similarity score (0-100) without any additional explanation
    or justification.
Template 16: Prompt generated by Gemini for Llama 3 70B evaluation.
SYSTEM: You are a medical diagnosis evaluator. Your task is to compare two medical text
summaries and assess the degree to which the first text (the "Candidate Answer") captures the
essential diagnostic information contained in the second text (the "Reference Answer"). Focus
on the core medical meaning and the completeness of the diagnoses, not on the writing style or
specific wording.
USER: INSTRUCTIONS:
1. Read the "Reference Answer" carefully. Identify the key diagnoses (both primary and
differential) and any crucial supporting information (e.g., symptoms, test results, ruled-out
disorderss).
2. Read the "Candidate Answer". The "Candidate Answer" will be a concise list of diagnoses and
key findings.
3. Compare the "Candidate Answer" to the "Reference Answer" based on the following criteria:
    • Diagnostic Accuracy: Does the "Candidate Answer" correctly identify the most likely
    diagnosis (if one is clearly indicated in the "Reference Answer")?
    • Diagnostic Completeness: Does the "Candidate Answer" include all of the major
    differential diagnoses mentioned in the "Reference Answer"? It is acceptable if the
    "Candidate Answer" uses slightly different terminology as long as the underlying medical
    concept is the same.
    • Supporting Information: Does the "Candidate Answer" include the key symptoms, test
    results, or other evidence that support the diagnoses? It doesn’t need to include every
    detail, but it should capture the most important ones. It’s ok if this supporting
    information is implicitly present in the reference answer.
    • Exclusion of Irrelevant Information: Do not penalize the "Candidate Answer" for including
    additional information that is not present in the "Reference Answer", as long as that
    information is medically accurate and relevant to the case.
    • groundance does not affect the score: Do not penalize a response for being redundant.
4. Provide a single numerical similarity score between 0 and 100, where:
    • 0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms
    of diagnosis or supporting information. It’s completely wrong.
    • 50: The "Candidate Answer" identifies some relevant diagnoses or supporting information,
    but it misses major elements or includes significant inaccuracies.
    • 75: The "Candidate Answer" correctly identifies the primary diagnosis (if applicable) and
    most of the important differential diagnoses and supporting information. It may miss some
    minor details or include some slight inaccuracies.
    • 90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core
    diagnostic meaning and completeness. It may have minor omissions or differences in
    emphasis, but it captures almost all of the essential information.
    • 100: The "Candidate Answer" is a perfect match in terms of core diagnostic meaning. It
    includes all of the key diagnoses and supporting information from the "Reference Answer",
    even if the wording or organization is different.
5. Do not justify or explain the answer. Give a score between 0 and 100, no other text.
Here is an example for better understanding:
Reference Answer:
Differential diagnosis: 1. Delirium 2. Substance intoxication 3.Substance withdrawal delirium
4. Mental retardation
Candidate Answer:
- Primary Diagnosis: Alzheimer’s_Disease
- Key Findings: Progressive memory loss, disorientation, difficulty with ADLs.
- Differential Diagnoses: Vascular_Dementia, Lewy_Body_Dementia, Frontotemporal_Dementia
ASSISTANT: 75
Here is another example:
Reference Answer:
Alcohol Use Disorder (Severe)
Candidate Answer:
- Primary Diagnosis: Alcohol_Use_Disorder (Mild)
- Key Findings: History of heavy alcohol use, withdrawal_symptoms.
ASSISTANT: 50
Now, evaluate the following:
Template 17: Treatment Prediction Prompt
You are a psychiatrist. Based on the patient’s clinical history, answer the following question
with a medically appropriate and individualized treatment recommendation.
Guidelines
- Base your treatment recommendation primarily on the Patient’s History. Use information from
the Mental and Physical Examinations only if it directly supports your treatment choice (if
found).
- Your response should be clinically sound, concise, and focused on the patient’s specific
presentation.
- Justify your recommendation using evidence from the case. Do not include general advice or
unrelated commentary.
- Respond in full, professional sentences.
Patient’s History :{History}
Mental State Examination : {Mental State Examination}
Physical Examination : {Physical Examination}
Question : {Question}
Template 18: Treatment Evaluation Prompt
SYSTEM: You are a medical treatment evaluator. Your task is to compare two medical treatment
summaries and assess the degree to which the first text (the "Candidate Answer") captures the
essential therapeutic interventions described in the second text (the "Reference Answer").
Focus on the medical appropriateness and completeness of the treatment plan, *not on the
writing style or specific wording.
USER:
INSTRUCTIONS:
1. Read the "Reference Answer" carefully. Identify the key components of the treatment plan,
including primary therapeutic interventions, medications, behavioral strategies, or other
relevant clinical actions.
2. Read the "Candidate Answer". The "Candidate Answer" will be a concise summary of proposed
treatments and strategies.
3. Compare the "Candidate Answer" to the "Reference Answer" based on the following criteria:
     Treatment Accuracy: Does the "Candidate Answer" correctly identify the key treatment or
    intervention as outlined in the "Reference Answer"?
     Treatment Completeness: Does the "Candidate Answer" include *all of the major treatment
    components mentioned in the "Reference Answer"? Slightly different terminology is
    acceptable as long as the clinical *concept is preserved.
     Supporting Rationale: Does the "Candidate Answer" capture the reasoning or context for
    the treatment (e.g., why a specific method or medication is chosen)? It doesn’t need to
    list every detail, but should reflect the main justification.
     Exclusion of Irrelevant Information: Do *not penalize the "Candidate Answer" for
    including additional, medically accurate and relevant information not found in the
    "Reference Answer".
     groundance does not affect the score: Do not penalize a response for being redundant.
4. Provide a single numerical similarity score between 0 and 100, where:
     0: The "Candidate Answer" has no meaningful overlap with the "Reference Answer" in terms
    of treatment or rationale. It is completely inappropriate or unrelated.
     50: The "Candidate Answer" identifies *some relevant treatment components or
    strategies, but misses key parts or includes significant inaccuracies.
     75: The "Candidate Answer" correctly identifies the primary treatment and most important
    additional strategies or reasoning. Minor omissions or slight inaccuracies are acceptable.
     90: The "Candidate Answer" is very close to the "Reference Answer" in terms of core
    treatment strategies and rationale. It may have minor differences in emphasis or detail.
     100: The "Candidate Answer" is a perfect match in terms of treatment content and
    rationale. It includes all key interventions and supporting reasoning from the "Reference
    Answer", even if the wording or structure is different.
5. Do not justify or explain the answer. Give a score between 0 and 100, no other text.
Here is an example for better understanding:
Reference Answer:
Treatment includes structured desensitization therapy and cognitive behavioral therapy (CBT).
Desensitization is done through exposure to a graded hierarchy of fear situations, supported
by relaxation techniques and possibly hypnotherapy. Treatment of the current abscess under
sedation is required first, followed by referral to a specialist dentist and CBT led by a
mental health professional.
Candidate Answer:
- Suggested Treatment: CBT for dental phobia
- Steps: Desensitization with imagery and real-life exposure, use of sedation for abscess
treatment, follow-up with a specialist dentist
ASSISTANT: 100
Here is another example:
Reference Answer:
CBT and desensitization therapy are the main treatments. Use of sedation is recommended
initially to manage acute dental issues. Follow-up should include exposure-based therapy and
psychoeducation.
Candidate Answer:
- Suggested Treatment: Use of medication for phobia and pain
- Notes: Patient should be given anti-anxiety medications and then referred to a dentist
ASSISTANT: 50
Now, evaluate the following:
Reference Answer:
{Answer}
Candidate Answer:
{response}
10.5 Classification Mapping
Template 19: Specific Disorders List
[ "Intellectual Developmental Disorder (Intellectual disability)",
"Global Developmental Delay", "Unspecified Intellectual Developmental Disorder",
"Language Disorder", "Speech Sound Disorder", "Childhood-Onset Fluency Disorder (Stuttering)",
"Social (Pragmatic) Communication Disorder", "Unspecified Communication Disorder",
"Autism Spectrum Disorder", "Attention-Deficit/Hyperactivity Disorder",
"Other Specified Attention-Deficit/Hyperactivity Disorder",
"Unspecified Attention-Deficit/Hyperactivity Disorder","Specific Learning Disorder",
"Developmental Coordination Disorder", "Stereotypic Movement Disorder",
"Substance-Induced Repetitive Behaviors", "Tourette’s Disorder",
"Persistent (Chronic) Motor or Vocal Tic Disorder", "Provisional Tic Disorder",
"Other specified Tic disorder", "Unspecified Tic Disorder",
"Other specified Neurodevelopmental disorders",
"Unspecified neurodevelopmental disorders","Delusional Disorder", "Brief Psychotic Disorder",
"Schizophreniform Disorder", "Schizophrenia", "Schizoaffective Disorder",
"Substance/Medication-Induced Psychotic Disorder",
"Psychotic disorder due to another medical disorders",
"Catatonia associated with another mental disorder",
"Catatonia associated with another medical disorder", "Unspecified Catatonia",
"Other Specified Schizophrenia Spectrum and Other Psychotic Disorder",
"Unspecified Schizophrenia Spectrum and Other Psychotic Disorder", "Bipolar I Disorder",
"Bipolar II Disorder", "Cyclothymic Disorder",
"Substance/Medication-Induced Bipolar and Related Disorder",
"Bipolar and Related Disorder Due to Another Medical disorders",
"Other Specified Bipolar and Related Disorder", "Unspecified Bipolar and Related Disorder",
"Unspecified Mood Disorder", "Disruptive Mood Dysregulation Disorder",
"Major Depressive Disorder", "Persistent Depressive Disorder",
"Premenstrual Dysphoric Disorder", "Substance/Medication-Induced Depressive Disorder",
"Depressive Disorder Due to Another Medical disorders",
"Other Specified Depressive Disorder", "Unspecified Depressive Disorder",
"Separation Anxiety Disorder", "Selective Mutism", "Specific Phobia","Social Anxiety Disorder",
"Panic Disorder", "Agoraphobia","Generalized Anxiety Disorder",
"Substance/Medication-Induced Anxiety Disorder",
"Anxiety Disorder Due to Another Medical disorders", "Other Specified Anxiety Disorder",
"Unspecified Anxiety Disorder", "Obsessive-Compulsive Disorder", "Body Dysmorphic Disorder",
"Hoarding Disorder","Trichotillomania (Hair-Pulling Disorder)",
"Excoriation (Skin-Picking) Disorder",
"Substance/Medication-Induced Obsessive-Compulsive and Related Disorder",
"Obsessive-Compulsive and Related Disorder Due to Another Medical disorders",
"Other Specified Obsessive-Compulsive and Related Disorder",
"Unspecified Obsessive-Compulsive and Related Disorder", "Reactive Attachment Disorder",
"Disinhibited Social Engagement Disorder", "Posttraumatic Stress Disorder",
"Acute Stress Disorder", "Adjustment Disorders", "Prolonged Grief Disorder",
"Other Specified Trauma- and Stressor-Related Disorder",
"Unspecified Trauma- and Stressor-Related Disorder", "Dissociative Identity Disorder",
"Dissociative Amnesia", "Depersonalization/Derealization Disorder",
"Other Specified Dissociative Disorder", "Unspecified Dissociative Disorder",
"Somatic Symptom Disorder", "Illness Anxiety Disorder",
"Functional Neurological Symptom Disorder (Conversion Disorder)",
"Psychological Factors Affecting Other Medical disorderss", "Factitious Disorder",
"Other Specified Somatic Symptom and Related Disorders",
"Unspecified Somatic Symptom and Related Disorder", "Pica", "Rumination Disorder",
"Avoidant/Restrictive Food Intake Disorder", "Anorexia Nervosa", "Bulimia Nervosa",
"Binge-eating disorder", "Other Specified Feeding or eating disorder",
"Unspecified Feeding or eating disorder", "Enuresis", "Encopresis",
"Other Specified Elimination Disorder", "Unspecified Elimination Disorder",
"Insomnia Disorder", "Hypersomnolence Disorder", "Narcolepsy",
"Obstructive Sleep Apnea Hypopnea", "Central Sleep Apnea",
"Sleep-Related Hypoventilation", "Delayed Sleep Phase Type",
"Advanced Sleep Phase Type", "Irregular Sleep-Wake Type",
"Non-24-Hour Sleep-Wake Type", "Shift Work Type", "Parasomnias",
"Non-Rapid Eye Movement Sleep Arousal Disorders", "Nightmare Disorder",
"Rapid Eye Movement Sleep Behavior Disorder", "Restless Legs Syndrome",
"Substance/Medication-Induced Sleep Disorder", "Other Specified Insomnia Disorder",
"Unspecified Insomnia Disorder", "Other Specified Hypersomnolence Disorder",
"Unspecified Hypersomnolence Disorder", "Other Specified Sleep-Wake Disorder",
"Unspecified Sleep-Wake Disorder", "Delayed Ejaculation", "Erectile Disorder",
"Female Orgasmic Disorder", "Female Sexual Interest/Arousal Disorder",
"Genito-Pelvic Pain/Penetration Disorder", "Male Hypoactive Sexual Desire Disorder",
"Premature (Early) Ejaculation", "Substance/Medication-Induced Sexual Dysfunction",
"Other Specified Sexual Dysfunction", "Unspecified Sexual Dysfunction",
"Other Specified Gender Dysphoria", "Unspecified Gender Dysphoria",
"Oppositional Defiant Disorder","Intermittent Explosive Disorder","Conduct Disorder",
"Pyromania", "Kleptomania",
"Other Specified Disruptive, Impulse-Control, and Conduct Disorder",
"Unspecified Disruptive, Impulse-Control, and Conduct Disorder",
"Alcohol Use Disorder", "Alcohol Intoxication", "Alcohol Withdrawal",
"Alcohol-induced Mental disorders", "Unspecified Alcohol-Related Disorder",
"Caffeine Intoxication", "Caffeine Withdrawal", "Caffeine-Induced Mental Disorders",
"Unspecified Caffeine-Related Disorder", "Cannabis Use Disorder",
"Cannabis Intoxication", "Cannabis Withdrawal", "Cannabis-Induced Mental Disorders",
"Unspecified Cannabis-Related Disorder", "Phencyclidine Use Disorder",
"Other Hallucinogen Use Disorder", "Phencyclidine Intoxication",
"Other Hallucinogen Intoxication", "Hallucinogen Persisting Perception Disorder",
"Phencyclidine-Induced Mental Disorders", "Hallucinogen-Induced Mental Disorders",
"Unspecified Phencyclidine-Related Disorder",
"Unspecified Hallucinogen-Related Disorder", "Inhalant Use Disorder",
"Inhalant Intoxication", "Inhalant-Induced Mental Disorders",
"Unspecified Inhalant-Related Disorder","Opioid Use Disorder","Opioid Intoxication",
"Opioid Withdrawal", "Opioid-Induced Mental Disorders",
"Unspecified Opioid-Related Disorder",
"Sedative, Hypnotic, or Anxiolytic Use Disorder",
"Sedative, Hypnotic, or Anxiolytic Intoxication",
"Sedative, Hypnotic, or Anxiolytic Withdrawal",
"Sedative-, Hypnotic-, or Anxiolytic-Induced Mental Disorders",
"Unspecified Sedative-, Hypnotic-, or Anxiolytic- ]
Template 20: Disorder Categories List
["Intellectual Developmental Disorders","Communication Disorders",
"Autism Spectrum Disorder","Attention-Deficit/Hyperactivity Disorder",
"Specific Learning Disorder","Motor Disorders","Tic Disorders",
"Schizophrenia Spectrum and Other Psychotic Disorders",
"Bipolar and Related Disorders","Depressive Disorders","Anxiety Disorders",
"Obsessive-Compulsive and Related Disorders","Trauma- and Stressor-Related Disorders",
"Dissociative Disorders","Somatic Symptom and Related Disorders",
"Feeding and Eating Disorders","Elimination Disorders","Sleep-Wake Disorders",
"Breathing-Related Sleep Disorders","Parasomnias","Sexual Dysfunctions","Gender Dysphoria",
"Disruptive, Impulse-Control, and Conduct Disorders",
"Substance-Related and Addictive Disorders","Neurocognitive Disorders",
"Personality Disorders","Medication-induced Movement Disorders","Paraphilic Disorders"]